### 05 Aug 2021 v1

# Geometric Embeddability of Complexes is $\exists\mathbb{R}$-complete

Abstract
We show that the decision problem of determining whether a given (abstract simplicial) $k$-complex has a geometric embedding in $\mathbb{R}^{d}$ is complete for the Existential Theory of the Reals for all $d\geq 3$ and $k\in\{d-1,d\}$. This implies that the problem is polynomial time equivalent to determining whether a polynomial equation system has a real root. Moreover, this implies NP-hardness and constitutes the first hardness results for the algorithmic problem of geometric embedding (abstract simplicial) complexes.
Figure 1: Illustration of different embeddings of complexes; figure taken from Bing [8, Annals of Mathematics 19591959].
1 Introduction
Since the dawn of the last century, much attention has been devoted to studying embeddings of complexes [29, 8, 20, 27, 28, 37, 48, 59]. Typical types of embeddings include geometric (also referred to as linear), piecewise linear (PL), and topological embeddings, see also Figure 1. For formal definitions, we refer to Section 1.2; here we give an illustrative example. Embeddings of a $1$-complex in the plane correspond to drawings of a graph in the plane. In a topological embedding, each edge is represented by a Jordan arc, in a PL embedding it is a concatenation of a finite number of segments, and in a geometric embedding each edge is a segment. Unlike in higher dimensions, for the embeddability of complexes in the plane all three notions coincide.
We are interested in the problem of deciding whether a given $k$-complex has a linear/piecewise linear/topological embedding in $\mathbb{R}^{d}$. Several necessary and sufficient conditions are easy to identify and have been known for many decades. For instance, a $k$-simplex requires $(k+1)$ points in general position in $\mathbb{R}^{d}$ and, thus, $k\leq d$ is an obvious necessary condition. Moreover, it is straight-forward to verify that every set of $n$ points in $\mathbb{R}^{3}$ in general position allows for a geometric embedding of any 1-complex on $n$ vertices, i.e., straight-line drawing of a (complete) graph. Indeed, this fact generalizes to higher dimensions: every $k$-complex embeds (even linearly) in $\mathbb{R}^{2k+1}$ [37]. Van Kampen and Flores [24, 51, 59] showed that this bound is tight by providing $k$-complexes that do not topologically embed into $\mathbb{R}^{2k}$. For some time, it was believed that the existence of a topological embedding also implies the existence of a geometric embedding, e.g., Grünbaum conjectured that if a $k$-dimensional simplicial complex is topologically embeddable in $\mathbb{R}^{2k}$ then it also geometrically embeds in $\mathbb{R}^{2k}$ [27]. However, this was consequently disproven. In particular, for every $k\geq 2$ and $k+1\leq d\leq 2k$, there exist $k$-complexes that have a PL embedding, but no geometric embedding in $\mathbb{R}^{d}$ [9, 10]. In contrast, PL and topological embeddability coincides in many cases, e.g., if $d\leq 3$ [8, 43] or $d-k\geq 3$ [11]. There are many further necessary and sufficient conditions for geometric embeddings [6, 41, 42, 51, 56, 57] and PL or/and topological embeddings [19, 25, 44, 48, 58].
In recent years, the algorithmic complexity of deciding whether or not a given complex is embeddable gained attention. In the absence of a complete characterization, an efficient algorithm is the best tool to decide embeddability. For instance, embeddings of 1-complexes in the plane corresponds to testing graph planarity which is known to be polynomial time decidable [30]. Conversely, the proven non-existence of efficient algorithms may offer a rigorous proof that a complete characterization is impossible. To give a concrete example, let $\textsc{Embed}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a PL embedding in $\mathbb{R}^{d}$. Because $\textsc{Embed}_{4\rightarrow 5}$ is known to be undecidable [23, 53], we have a proof that there does not exist an efficient algorithm for $\textsc{Embed}_{4\rightarrow 5}$ – even without any complexity assumptions (such as $\mathrm{NP}\neq\mathrm{P}$ or similar).
More recently, there have been several breakthroughs concerning the PL-embeddability. For an overview of the state of the art, consider Table 1.
Table 1: Overview for PL
In dimensions $d\geq 4$, the decision problem $\textsc{Embed}_{k\rightarrow d}$ is polynomial-time decidable for $k<\nicefrac{{2}}{{3}}(d-1)$ [15, 12, 14, 31] and NP-hard for all remaining cases [35], i.e., for all $k\geq\nicefrac{{2}}{{3}}(d-1)$. For $d\leq 5$ and $k\in\{d,d-1\}$, $\textsc{Embed}_{k\rightarrow d}$ is even known to be undecidable; for all other NP-hard cases and $d\geq 4$ decidability is unknown. For the case $d=3$, Matoušek, Sedgwick, Tancer, and Wagner have shown that $\textsc{Embed}_{2\rightarrow 3}$ and $\textsc{Embed}_{3\rightarrow 3}$ are decidable [34] and de Mesmay, Rieck, Sedgwick, and Tancer have proved NP-hardness [38].
Building upon [35], Skopenkov and Tancer [55] proved NP-hardness for a relaxed notion called almost (PL/topological) embeddability where disjoint hyperedges must be mapped to disjoint objects. More precisely, they showed that recognizing almost embeddability of $k$-complexes in $\mathbb{R}^{d}$ is $\mathrm{NP}$-hard for all $d,k\geq 2$ such that $d\pmod{3}=1$ and $\nicefrac{{2}}{{3}}(d-1)\leq k\leq d$.
The analogous questions for geometric embeddings are wide open. Let $\textsc{GEM}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a geometric embedding in $\mathbb{R}^{d}$. In contrast to PL-embeddability, however, it is easy to see that $\textsc{GEM}_{k\rightarrow d}$} is decidable for all $k,d$. In analogy to the PL-embeddings, $\mathrm{NP}$-hardness has been conjectured by Skopenkov.
Conjecture ([54], Conjecture 3.2.2).
$\textsc{GEM}_{k\rightarrow d}$ is $\mathrm{NP}$-hard for all $k,d$ with
$\max\{3,k\}\leq d\leq\nicefrac{{3}}{{2}}k+1.$
Note that these pairs correspond to the NP-hard cases of $\textsc{Embed}_{k\rightarrow d}$.
Results.
In this work, we present the first results for any non-trivial entry with $d\geq 3$. More precisely, we establish the exact computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all values $d\geq 3$ and $k\in\{d,d-1\}$, hereby confirming the conjecture for these cases. This includes a complete understanding of the most intriguing entries with $d=3$. Table 2 summarizes the current knowledge on the computational complexity of $\textsc{GEM}_{k\rightarrow d}$.
Theorem 1.
For every $d\geq 3$ and every $k\in\{d,d-1\}$, the decision problem $\textsc{GEM}_{k\rightarrow d}$ is $\exists\mathbb{R}$-complete.
Moreover, the statement remains true even if a PL-embedding is given.
Table 2: Overview of the computational complexity of $\textsc{GEM}_{k\rightarrow d}$.
As $\mathrm{NP}\subseteq\exists\mathbb{R}$, our result confirms the conjecture by Skopenkov for the corresponding values of $k$ and $d$. Moreover, if $\mathrm{NP}\neq\exists\mathbb{R}$, $\textsc{GEM}_{k\rightarrow d}$ cannot be tackled with well developed tools for $\mathrm{NP}$-complete problems such as SAT solvers and ILPs. For more details, we refer to Section 1.1.
As a matter of fact, our proof implies that distinguishing between $k$-complexes with PL and geometric embeddings in $\mathbb{R}^{d}$ is complete for $\exists\mathbb{R}$.
Geometric embeddings of complexes can also be viewed as a representation of hypergraphs in which every edge is represented by a simplex. Of particular interest is the case of uniform hypergraphs where all hyperedges have the same number of elements. Thus, in the language of hypergraphs, our result reads as follows.
Corollary 2.
For all $d\geq 3$ and every $k\in\{d,d-1\}$, deciding whether a $(k+1)$-uniform hypergraph has a simplicial representation in $\mathbb{R}^{d}$ is $\exists\mathbb{R}$-complete.
Outline and Techniques.
The main idea to prove hardness of $\textsc{GEM}_{2\rightarrow 3}$} is to reduce from the Stretchability problem. In Stretchability, we are given a set of pseudolines (curves) in the plane and we are asked to decide if there are straight lines that intersect each other in the same combinatorial pattern as the pseudolines. See below for a formal definition. Given a pseudoline arrangement $L$, we construct a $2$-complex $C$, such that $C$ is realizable in $\mathbb{R}^{3}$ if and only if $L$ is stretchable.
In our construction each pseudoline is represented by a special edge. We wrap tunnels around the special edges and all tunnels are connected by triangles to a special apex vertex $u$. It will be straightforward to verify that if $L$ is stretchable, then the resulting simplex $C$ is embeddable. The other direction is more subtle: We show that if $C$ is embeddable, then seen from $u$, the special edges will appear as a realization of $L$ using straight lines, and hence $L$ is stretchable. The idea behind the proof is that the tunnels and the apex $u$ force together the correct combinatorics of the intersection pattern of the special edges seen from $u$.
In order to show hardness of $\textsc{GEM}_{3\rightarrow 3}$}, we use a similar construction, but where we “fatten” each triangle to a tetrahedron, by adding extra vertices.
We finally make a dimension reduction by proving that $\textsc{GEM}_{k\rightarrow d}$} reduces to $\textsc{GEM}_{k+1\rightarrow d+1}$}. Given a $k$-complex $C$, we create a $(k+1)$-complex $C^{+}$ in the following way. The vertices of $C^{+}$ are those of $C$ and two extra vertices $a$ and $b$. For each edge $e$ of $C$, we have the additional edges $e\cup\{a\}$ and $e\cup\{b\}$ in $C^{+}$. We prove that $C$ geometrically embeds in $\mathbb{R}^{d}$ if and only if $C^{+}$ geometrically embeds in $\mathbb{R}^{d+1}$.
Thus, we show that deciding whether a given complex with a PL-embedding also allows a geometric embedding is $\exists\mathbb{R}$-complete, i.e., distinguishing PL-embeddable and geometric embeddable complexes is $\exists\mathbb{R}$-complete.
1.1 Existential Theory of the Reals
The class of the existential theory of the reals $\exists\mathbb{R}$ (pronounced as ‘exists R’, ‘ER’, or ‘ETR’) is a complexity class which has gained a lot of interest in recent years, specifically in the computational geometry community. To define this class, we first consider the algorithmic problem Existential Theory of the Reals (ETR). An instance of this problem consists of a sentence of the form
$\exists x_{1},\ldots,x_{n}\in\mathbb{R}:\Phi(x_{1},\ldots,x_{n}),$
where $\Phi$ is a well-formed quantifier-free formula over the alphabet $\{0,1,+,\cdot,\geq,>,\wedge,\vee,\neg\}$, and the goal is to check whether this sentence is true. As an example of an ETR-instance, consider $\exists x,y\in\mathbb{R}\colon\Phi(x,y)=(x\cdot y^{2}+x\geq 0)\wedge\neg(y<2x)$, for which the goal is to determine whether there exist real numbers $x$ and $y$ satisfying the formula $\Phi(x,y)$.
The complexity class $\exists\mathbb{R}$ is the family of all problems that admit a polynomial-time many-one reduction to ETR. It is known that
$\textrm{NP}\subseteq\exists\mathbb{R}\subseteq\textrm{PSPACE}.$
The first inclusion follows from the definition of $\exists\mathbb{R}$. Showing the second inclusion was first established by Canny in his seminal paper [16]. The complexity class $\exists\mathbb{R}$ gains its significance because a number of well-studied problems from different areas of theoretical computer science have been shown to be complete for this class.
Famous examples from discrete geometry are the recognition of geometric structures, such as unit disk graphs [36], segment intersection graphs [33], stretchability [40, 50], and order type realizability [33]. Other $\exists\mathbb{R}$-complete problems are related to graph drawing [32], Nash-Equilibria [7, 26], geometric packing [5], the art gallery problem [3], non-negative matrix factorization [49], polytopes [21, 46], geometric linkage constructions [1], training neural networks [4], visibility graphs [18], continuous CSPs [39], and convex covers [2]. The fascination for the complexity class stems not merely from the number of $\exists\mathbb{R}$-complete problems but from the large scope of seemingly unrelated $\exists\mathbb{R}$-complete problems. We refer the reader to the lecture notes by Matoušek [33] and surveys by Schaefer [47] and Cardinal [17] for more information on the complexity class $\exists\mathbb{R}$.
1.2 Definitions
Simplex.
A $k$-simplex is a $k$-dimensional polytope which is the convex hull of its $k+1$ vertices. Hence, a 0-simplex corresponds to a point, a 1-simplex to a segment, and a 2-simplex to a triangle etc. The convex hull of any nonempty proper subset of the $k+1$ points that define a $k$-simplex is called a face of the simplex. A simplicial complex $K$ is a set of simplices satifying the following two conditions: (i) Every face of a simplex from $K$ is also in $K$. (ii) For any two simplices $\sigma_{1},\sigma_{2}\in K$ with a non-empty intersection, the intersection $\sigma_{1}\cap\sigma_{2}$ is a face of $\sigma_{1}$ and $\sigma_{2}$. The purely combinatorial counterpart to a simplicial complex is an abstract simplicial complex, which we refer to simply as a complex.
Complex.
A complex $C=(V,E)$ is a finite set $V$ together with a collection of subsets $E\subseteq 2^{V}$ which is closed under taking subsets, i.e., $e\in E$ and $e^{\prime}\subseteq e$ imply that $e^{\prime}\in E$. A $k$-complex is a complex where the largest subset contains exactly $(k+1)$ elements. We call a complex pure if all (inclusion-wise) maximal elements in $E$ have the same cardinality.
The neighbors of every vertex $v\in V$ in a $k$-complex $C=(V,E)$ define a lower dimensional complex $C_{v}:=(V^{\prime},E^{\prime})$ where $E^{\prime}:=\{e\setminus\{v\}|v\in e\in E\}$ and $V^{\prime}:=N(v)=\bigcup_{e\in E^{\prime}}e$ are the neighbors of $v$.
Complexes are in close relation to Hypergraphs.
Hypergraphs.
Hypergraphs generalize graphs by allowing edges to contain any number of vertices. Formally, an undirected hypergraph $H$ is a pair $H=(X,E)$ where $X$ is a set of vertices, and $E$ is a set of non-empty subsets of $X$ called hyperedges (or edges). A $k$-uniform hypergraph is a hypergraph such that all its hyperedges contain exactly $k$ elements. Note that the maximal sets of a pure $k$-complex yield a $(k+1)$-uniform hypergraph and vice versa. Hence, $(k+1)$-uniform hypergraphs and pure $k$-complexes are in a straight-forward one-to-one correspondence. A simplicial representation of a $(k+1)$-uniform hypergraph is a geometric embedding of the corresponding complex.
Geometric embeddings.
A geometric embedding of a complex $C=(V,E)$ in $\mathbb{R}^{d}$ is a function $\varphi\colon V\rightarrow\mathbb{R}^{d}$ fulfilling the following two properties: (i) for every $e\in E$, $\overline{\varphi}(e):={\textrm{conv}}(\{\varphi(v)\colon v\in e\})$ is a simplex of dimension |e|1|e|-1 and (ii) for every pair $e,e^{\prime}\in E$, it holds that
$\overline{\varphi}(e)\cap\overline{\varphi}(e^{\prime})=\overline{\varphi}(e\cap e^{\prime}).$
Note that if $\varphi$ is a geometric embedding, then $\{\overline{\varphi}(e):e\in E\}$ is a simplicial complex. The problem $\textsc{GEM}_{k\rightarrow d}$ asks whether a given $k$-complex has a geometric embedding in $\mathbb{R}^{d}$.
PL-embeddings and topological embeddings.
In contrast to geometric embeddings, for PL or topological embeddings it is not sufficient to describe the mapping of the vertices. Thus, we define the embedding of a complex by its corresponding simplicial complex.
A PL/topological embedding of a simplicial complex $C=(V,E)$ in $\mathbb{R}^{d}$ is a piecewise/continuous linear function $\varphi\colon E\rightarrow\mathbb{R}^{d}$ such that the following property is satisfied: for every pair $e,e^{\prime}\in E$, it holds that
$\varphi(e)\cap\varphi(e^{\prime})=\varphi(e\cap e^{\prime}).$
Stretchability.
A pseudoline arrangement is a family of curves that apart from the ‘straightness’ share similar properties with a line arrangement. More formally, we define a (Euclidean) pseudoline arrangement as an arrangement of $x$-monotone curves in the Euclidean plane such that every pair of pseudolines meets exactly once in a crossing point. In fact, each pseudoline arrangement can be nicely encoded by a wiring diagram; see also Figure 2(a). A pseudoline arrangement is stretchable if it is combinatorially equivalent to an arrangement of straight-lines, i.e., if both arrangements are isomorphic. Stretchability denotes the algorithmic problem of deciding whether a given pseudoline arrangement is stretchable. In a seminal paper, Shor [50] proved that Stretchability is complete for the existential theory of the reals; for a stream-line exposition see Matoušek [33].
2 The Proof
Our proof of Theorem 1 consists of three parts.
  1. a)
    We establish $\exists\mathbb{R}$-membership. (Lemma 3, Section 2.1)
  2. b)
    We show $\exists\mathbb{R}$-hardness in $\mathbb{R}^{3}$, i.e., of $\textsc{GEM}_{2\rightarrow 3}$ and $\textsc{GEM}_{3\rightarrow 3}$. (Lemmas 4 and 5, Section 2.2)
  3. c)
    We reduce $\textsc{GEM}_{k\rightarrow d}$ to $\textsc{GEM}_{k+1\rightarrow d+1}$. (Lemma 6, Section 2.3)
Together Lemmas 3, 4, 5 and 6 prove Theorem 1.
2.1 Membership
In this subsection, we show $\exists\mathbb{R}$-membership of $\textsc{GEM}_{k\rightarrow d}$. Note that this is essentially folklore [13]. We give here a proof for the sake of completeness.
Lemma 3.
For all $k,d\in\mathbb{N}$, the decision problem $\textsc{GEM}_{k\rightarrow d}$ is contained in $\exists\mathbb{R}$.
Proof.
In order to show membership in $\exists\mathbb{R}$, we use the following characterization by Erickson, Hoog and Miltzow [22]: A problem $P$ lies in $\exists\mathbb{R}$ if and only if there exists a real verification algorithm $A$ for $P$ that runs in polynomial time on the real RAM. In particular, for every yes-instance $I$ of $P$ there exists a polynomial sized witness $w$ such that $A(I,w)$ returns yes, and for every no-instance $I$ of $P$ and any witness $w$, $A(I,w)$ returns no. In contrast to the definition of the complexity class NP, we also allow witnesses that consist of real numbers. Consequently, we execute $A$ on the real RAM as well.
It remains to present a real verification algorithm for $\textsc{GEM}_{k\rightarrow d}$. While the witness describes the coordinates of the vertices, the algorithm checks for intersections between any two simplices. Note that each simplex is a convex set and the intersection is a convex set as well. If the simplex $S$ has $n$ vertices, then we can efficiently determine $n$ linear inequalities and at most one linear equality that together describe $S$. Now checking for intersections can be reduced to linear programming, which is polynomial time solvable. This finishes the description of the real verification algorithm. ∎
2.2 Hardness in three dimensions
This section is dedicated to proving Theorem 1 for $k=2,3$ and $d=3$. The crucial part lies in the case $k=2$.
Lemma 4.
The decision problem $\textsc{GEM}_{2\rightarrow 3}$ is $\exists\mathbb{R}$-hard.
Proof.
We reduce from the $\exists\mathbb{R}$-hard problem Stretchability, as described in Section 1.2. Let $L$ be an arrangement of $n$ pseudolines in the plane. Every pseudoline arrangement has a representation as a wiring diagram where each pseudoline is represented by a monotone curve consisting of 2n12n-1 segments; for an illustration consider Figure 2(a). We add a pseudoline at infinity $\ell_{0}$ that intersects all pseudolines in the beginning. We call the resulting pseudoline arrangement $L^{*}$. Note that $L^{*}$ is stretchable if and only if $L$ is stretchable. In the following, we construct a 2-complex $C=(V,E)$ that allows for a geometric realization if and only if $L^{*}$ (and thus $L$) is stretchable.
Construction of the 2-complex.
In order to define $C$, we use the intended geometric embedding. We will refer to the subsets in $C$ as vertices, edges, and triangles if they contain one, two or three elements, respectively. The construction has three steps.
Firstly, we place the pseudolines in 3-space such that each pseudoline $\ell_{i}$ lies the plane $z=i$. In particular, note that no two pseudolines intersect. Therefore, we can surround each (lifted) pseudoline by a triangulated tube; see also Figure 2(b).
(a) A pseudoline arrangement $L^{*}$.
(b) Construction of a tube.
(c) A section of a tube.
[…]
The tube is formed by $2n+1$ sections such that every second section covers a crossing (when viewed from the top). Moreover, each section consists of six triangles forming a triangulated triangular prism as illustrated in Figure 2(c). We close the tube with triangles at the ends and think of the bottom side of the prism to lie in the plane $z=i$. Hence, the tube has a distinguished top polyline.
Secondly, for pseudoline $\ell_{i}$, we insert a special edge $e_{i}$ between the two top vertices on either end of the tube; for later reference, we denote the end vertices by $s_{i}$ and $t_{i}$. The special edge $e_{i}$ is intended to lie inside the tube and represent the stretched pseudoline (if possible). To this end, we add a new vertex to the construction that we call apex and denote by $u$. Specifically, the edges of the triangles incident to $s_{i}$ and $t_{i}$ form a triangle with $u$; for an illustration, consider Figure 3.
(a) front view
(b) top view
(c) tent of $s_{i}$
Figure 3: Construction of the triangles incident to $u$ and edges ’close’ to $s_{i}$ (or $t_{i}$); orange edges form triangles with $u$.
Thirdly, triangles with the apex $u$ and almost every edge of the tubes will ensure the correct combinatorics of $L$. The idea is that the apex forms a triangle with every edge that it sees from the top (at infinity). Note that the apex does not see the bottom side of any tube and that an invisible top part of a tube corresponds to a crossing in $L$. Figure 4(a) illustrates the neighborhood of the apex at crossings.
(a) apex triangles at crossings
(b) Neighborhood complex $\overline{C}_{u}$ corresponds to a 3-connected planar graph.
Figure 4: Neighborhood of the apex (around crossings)
As of now, the neighborhood of the apex $u$ induces a plane 2-connected graph $C^{\prime}_{u}$. Lastly, we add one vertex in each face of $C^{\prime}_{u}$ which is not a triangle, see Figure 4(b); note that these faces are in bijection with the faces of $L^{*}$. We call the now constructed complex $\overline{C}$. Consequently, the neighborhood $N(u)$ induces a $1$-complex $\overline{C}_{u}$ that corresponds to a $3$-connected planar graph. Our final complex $C$ consist of two copies of $\overline{C}$ in which the apex vertices are identified. (We will later use these two copies in order to guarantee that the apex lies in the unbounded region for one copy of $\overline{C}$.)
Time Complexity.
In order to verify that the construction shows $\exists\mathbb{R}$-hardness, we argue that it has a polynomial running time in the size of the input. To this end, note that a pseudoline arrangement with $n$ pseudolines can be described by the $O(n^{2})$ crossings. Thus,the input size is $N=O(n^{2})$. It is easy to see that our construction has a size linear in $N$: For each segment and crossing of the pseudoline arrangement, we build a constant size construction involving the apex. Consequently, the total construction has size $O(n^{2})=O(N)$.
Correctness.
It remains to show that $L$ is stretchable if and only if $C$ has a geometric embedding in $\mathbb{R}^{3}$.
If $L$ is stretchable, the geometric embedding of $C$ is straight-forward and goes along the same lines as its construction: Considering the stretched line arrangement $L^{\prime}$ equivalent to $L$, we add $\ell_{0}$ and lift the lines in 3-space, construct the tunnels, make a second copy, and can insert the apex $u$ (high enough) above. This yields a geometric embedding of $C$.
The reverse direction is a little more involved. Let $\varphi$ denote a geometric embedding of $C$. Note that the tunnel of each pseudoline forms a closed topological sphere; all of which are pairwise disjoint. By a generalization of the Jordan curve theorem (also known as the Jordan–Brouwer separation theorem) any topological embedding of a (d1)(d-1)-sphere in $\mathbb{R}^{d}$ splits the space into two components; we refer to the bounded component as its inside and to the unbounded as its outside. Note that the apex $u$ can lie inside at most one tunnel; otherwise, among all tunnels containing $u$, the innermost separates $u$ from the outermost. Consequently, for at least one one copy of $\overline{C}$ contained in $C$, the apex $u$ lies in the outside of each tunnel in $\overline{C}$. From now on, we focus on the geometric embedding of this $\overline{C}$ and do not make further use of the other copy.
As a crucial property, we argue that the special edge of each pseudoline lies in the interior of its tunnel. To this end, consider $s_{i}$ and the tunnel forming a closed topological sphere incident to $s_{i}$. Note that the triangles formed by the apex $u$ and objects of the last section of the tunnel (non-incident to $s_{i}$) form a tent on top of the tunnel-sphere, see Figure 3(c). In particular, the tent together with the tunnel triangles incident to $s_{i}$ form another sphere incident to $s_{i}$. If $e_{i}$ would start towards the outside of the tunnel, then it would be trapped inside the tent-sphere which is not incident to $t_{i}$. Consequently, $e_{i}$ lies inside the tunnel. Moreover, the edges $\{u,s_{i}\}$ and $\{u,t_{i}\}$ ensure that $e_{i}$ essentially stretches over the entire tunnel. Most importantly, $e_{i}$ runs through every section of the tunnel that belongs to a crossing.
Figure 5: Obtaining a line arrangement from a geometric embedding.
Now, we consider a small ball around the apex which does not contain any other vertex, see Figure 5(a). For each triangle containing the apex $u$, the intersection with the sphere $S$ yields an arc of a great circle. Consequently, we obtain a crossing-free drawing of the neighborhood complex $\overline{C}_{u}$ of $u$ with arcs of great circles on $S$. Because $\overline{C}_{u}$ is 3-connected, there exists a unique planar drawing on $S$. Recall that $\overline{C}_{u}$ captures the crossings of the tunnels and thus the combinatorics of $L^{*}$. Moreover, for each special edge $e_{i}$, we consider the projection $a_{i}$ of the (artificial) triangle $\{s_{i},t_{i},u\}$ onto the sphere $S$; note that $\{s_{i},t_{i},u\}$ is not a triangle of $C$, see Figure 5(b). The resulting arc $a_{i}$ lies within projection of the tunnel because $e_{i}$ is contained in the tunnel. Hence, the arc arrangement has the same combinatorics/intersection pattern as $L$, i.e., the order of intersections along each arc/pseudoline are the same.
Finally, we consider the supporting planes of the arcs (through $u$). Specifically, we consider the plane $p$ through the arc $a_{0}$ and apex $u$ and assume without loss of generality that $p$ is horizontal. Clearly, all arcs start to the same hemisphere of $S$ cut by $p$; let suppose the bottom hemisphere. Then, the intersection of any two arcs $a_{i}$ and $a_{j}$ lie on the bottom hemisphere; otherwise the arcs contain both intersections of the supporting great circles. Let $p^{\prime}$ denote the plane obtained by shifting $p$ slightly to the bottom. Consequently, we can ’project’ the arc arrangement onto the plane $p^{\prime}$ by considering the intersection of $p^{\prime}$ with the plane supporting the arc $a_{i}$. Clearly, this intersection yields a line. Hence, we obtain a line arrangement in the plane $p^{\prime}$ that proves the stretchability of $L$. ∎
Fattening the Complex.
The constructed 2-complex in the proof of Lemma 4 was not pure. Specifically, the special edges are not contained in any triangle. We can obtain a pure 2-complex by adding one new vertex to each special edge such that it forms a special triangle. Similarly, we can add a private vertex to each triangle to form a pure 3-complex $C^{\prime}$. Given a geometric embedding of $C$, these new vertices can easily be added close enough to their defining set in $C$. Hence, $C$ has a geometric embedding if and only if $C^{\prime}$ has a geometric embedding in $\mathbb{R}^{3}$. Together with Lemma 4, this very small modification shows hardness of $\textsc{GEM}_{3\rightarrow 3}$.
Lemma 5.
The decision problem $\textsc{GEM}_{3\rightarrow 3}$ is $\exists\mathbb{R}$-hard.
2.3 Dimension Reduction
In order to show hardness for all remaining cases of Theorem 1, we establish the following dimension reduction. For interesting dimension reduction in the context of PL-embeddings, we refer to [45, 52, 23, 44].
Lemma 6.
The decision problem $\textsc{GEM}_{k\rightarrow d}$ reduces to $\textsc{GEM}_{k+1\rightarrow d+1}$.
The idea is to add two apices to a $k$-complex $C$ in order to obtain a $(k+1)$-complex $C^{+}$. We will then argue that $C$ has a geometric embedding in $\mathbb{R}^{d}$ if and only if $C^{+}$ has a geometric embedding in $\mathbb{R}^{d+1}$. More formally, for a complex $C=(V,E)$ and a disjoint vertex set $U$, CUC*U denotes the join complex $(V\cup U,E^{\prime})$ where $E^{\prime}:=\{e\cup u|e\in E,u\in U\}$. The following claim immediately implies Lemma 6.
Claim 7.
Let $C=(V,E)$ be a complex, $a,b\notin V$ two new vertices, and $C^{+}:=C*\{a,b\}$ their join complex. Then $C$ has a geometric embedding in $\mathbb{R}^{d}$ if and only if $C^{+}$ has a geometric embedding in $\mathbb{R}^{d+1}$.
Proof.
Let $\varphi$ be a geometric embedding of $C$ in $\mathbb{R}^{d}$. Then, we define for $v\in V\cup\{a,b\}$,
$\varphi^{\prime}(v)=\begin{cases}(\ \ \varphi(v)\ \ ,\ 0\ )&\text{ if }v\in V,\\ (0,\dots,0,+1)&\text{ if }v=a,\\ (0,\dots,0,-1)&\text{ if }v=b.\end{cases}$
It is easy to check that $\varphi^{\prime}$ is a geometric embedding of $C^{+}$ in $\mathbb{R}^{d+1}$: While $a$ and $b$ are well separated in the last coordinate, all other potential intersections happen in the $d$-dimensional subspace induced by the first $d$ coordinates. Hence $\varphi$ implies the correctness of the geometric embedding.
For the reverse direction, consider a geometric embedding $\varphi$ of $C^{+}$ in $\mathbb{R}^{d+1}$. Let $\varphi_{a}:=\varphi(a)$ and $\varphi_{b}:=\varphi(b)$. Without loss of generality, we assume that $\varphi_{a}-\varphi_{b}$ is orthogonal to the first $d$ coordinates, i.e., $\varphi_{a}-\varphi_{b}$ is parallel to the $(d+1)$-st coordinate axis. Let $\overline{\varphi}(C):=\bigcup_{e\in E}\overline{\varphi}(e)$ denote the induced geometric subrepresenation of $C$. We claim that the function $f:\overline{\varphi}(C)\to\mathbb{R}^{d}$ obtained by restricting to the first $d$ coordinates is injective. Thus $\varphi^{\prime}:=f\circ\varphi$ yields a representation of $C$ in $\mathbb{R}^{d}$.
For the purpose of a contradiction, suppose that $f$ is not injective. Then there exist two distinct points $p=(p_{1},\ldots,p_{d+1})$ and $q=(q_{1},\ldots,q_{d+1})$ with $p,q\in\overline{\varphi}(C)$ such that $(p_{1},\ldots,p_{d})=(q_{1},\ldots,q_{d})$ and $p_{d+1}\neq q_{d+1}$. Without loss of generality, we may assume that $p_{d+1}>q_{d+1}$. Consider the plane $P$ spanned by $\varphi_{a},\varphi_{b},p$. Note that $q\in P$. For an illustration, see Figure 6.
Figure 6: The geometric embedding of $\varphi$ of $C^{+}$, gives a monotone embedding of $C$, otherwise we can find an intersection.
Let us denote with $e_{p}\in E$ and $e_{q}\in E$ any choice of hyperedges such that $p\in\overline{\varphi}(e_{p})$ and $q\in\overline{\varphi}(e_{q})$. Consider the two open segments ${\textrm{seg}}^{\circ}(\varphi_{a},q)\in\overline{\varphi}(e_{q}\cup a)$ and ${\textrm{seg}}^{\circ}(\varphi_{b},p)\in\overline{\varphi}(e_{p}\cup b)$. Clearly, these open segments intersect in a point $x$, as illustrated Figure 6. Because $\varphi$ is a geometric embedding, it holds that
$x\in\overline{\varphi}(e_{q}\cup a)\cap\overline{\varphi}(e_{p}\cup b)=\overline{\varphi}(e_{q}\cap e_{p})=\overline{\varphi}(e_{q})\cap\overline{\varphi}(e_{p}).$
[…]
3 Conclusion
We established the computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all $d\geq 3$ and $k\in\{d-1,d\}$. In particular, we showed that for these values it is complete for $\exists\mathbb{R}$ to distinguish PL-embeddable $k$-complexes in $\mathbb{R}^{d}$ from geometrically embeddable ones. Arguable, $\textsc{GEM}_{2\rightarrow 3}$ is the most interesting case.
Investigating the computational complexity for the remaining open entries in Table 2 remains for future work. We strengthen the conjecture of Skopenkov [54] as follows.
Conjecture.
The problem $\textsc{GEM}_{k\rightarrow d}$ is $\exists\mathbb{R}$-complete for all $k,d$ such that $\max\{3,k\}\leq d\leq 2k$.
Acknowledgements.
We thank Arkadiy Skopenkov for his kind and swift help with acquiring literature. Tillmann Miltzow is generously supported by the Netherlands Organisation for Scientific Research (NWO) under project no. 016.Veni.192.250. Linda Kleist was supported by a postdoc fellowship of the German Academic Exchange Service (DAAD).