### 05 Nov 2021 v2

# Geometric Embeddability of Complexes is $\exists\mathbb{R}$-complete

Abstract
We show that the decision problem of determining whether a given (abstract simplicial) $k$-complex has a geometric embedding in $\mathbb{R}^{d}$ is complete for the Existential Theory of the Reals for all $d\geq 3$ and $k\in\{d-1,d\}$. This implies that the problem is polynomial time equivalent to determining whether a polynomial equation system has a real solution. Moreover, this implies NP-hardness and constitutes the first hardness result for the algorithmic problem of geometric embedding (abstract simplicial) complexes.
Figure 1: Illustration of different embeddings of a complex; figure taken from Bing [undefg, Annals of Mathematics 19591959].
1 Introduction
Since the dawn of the last century, much attention has been devoted to studying embeddings of complexes [undefac, undefg, undeft, undefaa, undefab, undefal, undefay, undefaak]. Typical types of embeddings include geometric (also referred to as linear), piecewise linear (PL), and topological embeddings, see also Figure 1. For formal definitions, we refer to Section 1.2; here we give an illustrative example. Embeddings of a $1$-complex in the plane correspond to drawings of a graph in the plane. In a topological embedding, each edge is represented by a Jordan arc, in a PL embedding it is a concatenation of a finite number of segments, and in a geometric embedding each edge is represented by a segment. Unlike in higher dimensions, for the embeddability of complexes in the plane all three notions coincide.
We are interested in the problem of deciding whether a given $k$-complex has a linear/piecewise linear/topological embedding in $\mathbb{R}^{d}$. Several necessary and sufficient conditions are easy to identify and have been known for many decades. For instance, a $k$-simplex requires $k+1$ points in general position in $\mathbb{R}^{d}$ and, thus, $k\leq d$ is an obvious necessary condition. Moreover, it is straight-forward to verify that every set of $n$ points in $\mathbb{R}^{3}$ in general position allows for a geometric embedding of any 1-complex on $n$ vertices, i.e., the points are the vertices of a straight-line drawing of a (complete) graph. Indeed, this fact generalizes to higher dimensions: every $k$-complex embeds (even linearly) in $\mathbb{R}^{2k+1}$ [undefal]. Van Kampen and Flores [undefx, undefaae, undefaak] showed that this bound is tight by providing $k$-complexes that do not topologically embed into $\mathbb{R}^{2k}$. For some time, it was believed that the existence of a topological embedding also implies the existence of a geometric embedding, e.g., Grünbaum conjectured that if a $k$-complex topologically embeds in $\mathbb{R}^{2k}$, then it also geometrically embeds in $\mathbb{R}^{2k}$ [undefaa]. However, this was later disproven. In particular, for every $k,d\geq 2$ with $k+1\leq d\leq 2k$, there exist $k$-complexes that have a PL embedding in $\mathbb{R}^{d}$, but no geometric embedding in $\mathbb{R}^{d}$ [undefj, undefh, undefi]. In contrast, PL and topological embeddability coincides in many cases, e.g., if $d\leq 3$ [undefg, undefar] or $d-k\geq 3$ [undefk]. There are many further necessary and sufficient conditions known for geometric embeddings [undefe, undefap, undefaq, undefaae, undefaah, undefaai] and PL or/and topological embeddings [undefs, undefy, undefas, undefay, undefaaj, undefaag].
In recent years, the algorithmic complexity of deciding whether or not a given complex is embeddable gained attention. In the absence of a complete characterization, an efficient algorithm is the best tool to decide embeddability. For instance, deciding whether a 1-complex embeds in the plane corresponds to testing graph planarity and is thus polynomial time decidable [undefad]. Conversely, the proven non-existence of efficient algorithms may offer a rigorous proof that a complete characterization is impossible. To give a concrete example, let $\textsc{Embed}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a PL embedding in $\mathbb{R}^{d}$. Because $\textsc{Embed}_{4\rightarrow 5}$ is known to be undecidable [undefw, undefaac], we have a proof that there does not exist an efficient algorithm for $\textsc{Embed}_{4\rightarrow 5}$ – even without any complexity assumptions (such as $\mathrm{NP}\neq\mathrm{P}$ or similar).
More recently, there have been several breakthroughs concerning the PL embeddability. For an overview of the state of the art, consider Table 1.
Table 1: Overview of the complexity of $\textsc{Embed}_{k\rightarrow d}$.
In dimensions $d\geq 4$, the decision problem $\textsc{Embed}_{k\rightarrow d}$ is polynomial-time decidable for $k<\nicefrac{{2}}{{3}}\cdot(d-1)$ [undefo, undefl, undefn, undefaf] and NP-hard for all remaining cases [undefaj], i.e., for all $k\geq\nicefrac{{2}}{{3}}\cdot(d-1)$. For $d\geq 5$ and $k\in\{d-1,d\}$, $\textsc{Embed}_{k\rightarrow d}$ is even known to be undecidable; for all other NP-hard cases and $d\geq 4$ decidability is unknown. For the case $d=3$, Matoušek, Sedgwick, Tancer, and Wagner have shown that $\textsc{Embed}_{2\rightarrow 3}$ and $\textsc{Embed}_{3\rightarrow 3}$ are decidable [undefai] and de Mesmay, Rieck, Sedgwick, and Tancer have proved NP-hardness [undefam].
Building upon [undefaj], Skopenkov and Tancer [undefaaf] proved NP-hardness for a relaxed notion called almost (PL/topological) embeddability where it is only required that disjoint sets must be mapped to disjoint objects, i.e., this notions allows that two edges incident to a common vertex cross in an interior point. More precisely, they showed that recognizing almost embeddability of $k$-complexes in $\mathbb{R}^{d}$ is $\mathrm{NP}$-hard for all $d,k\geq 2$ such that $d\pmod{3}=1$ and $\nicefrac{{2}}{{3}}\cdot(d-1)\leq k\leq d$.
The analogous questions for geometric embeddings are wide open. Let $\textsc{GEM}_{k\rightarrow d}$ denote the algorithmic problem of determining whether a given $k$-complex has a geometric embedding in $\mathbb{R}^{d}$. In contrast to PL embeddability, however, it is easy to see that $\textsc{GEM}_{k\rightarrow d}$} is decidable for all $k,d$, since every instance can be expressed as a sentence in the first order theory of the reals, which is decidable. In analogy to the PL embeddings, $\mathrm{NP}$-hardness has been conjectured by Skopenkov.
Conjecture ([undefaad], Conjecture 3.2.2).
$\textsc{GEM}_{k\rightarrow d}$ is $\mathrm{NP}$-hard for all $k,d$ with
$\max\{3,k\}\leq d\leq\nicefrac{{3}}{{2}}\cdot k+1.$
Note that these parameters correspond to the NP-hard cases of $\textsc{Embed}_{k\rightarrow d}$, see also Table 1. Cardinal [undefq, Section 4] mentions $\textsc{GEM}_{2\rightarrow 3}$} as an interesting open problem. The closely related question of polyhedral complexes, posed in the Handbook of Discrete and Computational Geometry, reads as follows: When is a given finite poset isomorphic to the face poset of some polyhedral complex in a given space $\mathbb{R}^{d}$[undefax, Problem 20.1.1]. Note that simplicial complexes are special cases of polyhedral complexes, because each simplex is a basic polyhedron. The recognition of polyhedral complexes (with triangles and quadrangles) in $\mathbb{R}^{3}$ has been claimed to be $\exists\mathbb{R}$-complete [undefq, Theorem 5]. Focussing on convex polytopes, Richter-Gebert proved that recognizing convex polytopes in $\mathbb{R}^{4}$ is $\exists\mathbb{R}$-complete [undefau, undefav].
Our Results.
In this work, we present the first results concerning Skopenkov’s conjecture for any non-trivial entry with $d\geq 3$. More precisely, we establish the exact computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all values $d\geq 3$ and $k\in\{d-1,d\}$, hereby confirming the conjecture for these cases. This includes a complete understanding of the most intriguing entries with $d=3$. Note that this also answers the computational aspects of the question from the Handbook of Discrete and Computational Geometry. Table 2 summarizes the current knowledge on the computational complexity of $\textsc{GEM}_{k\rightarrow d}$.
Theorem 1.
For every $d\geq 3$ and each $k\in\{d-1,d\}$, the decision problem $\textsc{GEM}_{k\rightarrow d}$ is $\exists\mathbb{R}$-complete. Moreover, the statement remains true even if a PL embedding is given.
Table 2: Overview of the computational complexity of $\textsc{GEM}_{k\rightarrow d}$.
Our proof implies that distinguishing between $k$-complexes with PL and geometric embeddings in $\mathbb{R}^{d}$ is complete for $\exists\mathbb{R}$. Because $\mathrm{NP}\subseteq\exists\mathbb{R}$, our result confirms the conjecture by Skopenkov for the corresponding values of $k$ and $d$. Moreover, if $\mathrm{NP}\neq\exists\mathbb{R}$, the problem $\textsc{GEM}_{k\rightarrow d}$ cannot be tackled with well developed tools for $\mathrm{NP}$-complete problems such as SAT and ILP solvers. For more details, we refer to Section 1.1.
A geometric embedding of a complex can also be viewed as a simplicial representation of a hypergraph, i.e., a representation of a hypergraph in which every hyperedge is represented by a simplex. Of particular interest is the case of uniform hypergraphs where all hyperedges have the same number of elements. Thus, in the language of hypergraphs, our result reads as follows.
Corollary 2.
For all $d\geq 3$ and every $k\in\{d-1,d\}$, deciding whether a $(k+1)$-uniform hypergraph has a simplicial representation in $\mathbb{R}^{d}$ is $\exists\mathbb{R}$-complete.
Outline and techniques.
Our proof of Theorem 1 consists of three steps: Establishing $\exists\mathbb{R}$-membership, showing $\exists\mathbb{R}$-hardness in $\mathbb{R}^{3}$, i.e., of $\textsc{GEM}_{2\rightarrow 3}$ and $\textsc{GEM}_{3\rightarrow 3}$, and reducing $\textsc{GEM}_{k\rightarrow d}$ to $\textsc{GEM}_{k+1\rightarrow d+1}$. The core of the proof lies in establishing hardness of $\textsc{GEM}_{2\rightarrow 3}$}.
The main idea to prove hardness of $\textsc{GEM}_{2\rightarrow 3}$} is to reduce from the problem Stretchability. In Stretchability, we are given an arrangement of pseudolines (curves) in the plane and we are asked to decide whether there exists a set of straight lines that has the same combinatorial pattern as the pseudoline arrangement, see Figure 2(a) for an illustration and Section 1.2 for a formal definition. Given a pseudoline arrangement $L$, we construct a $2$-complex $C$ which has a geometric embedding in $\mathbb{R}^{3}$ if and only if $L$ is stretchable.
Figure 2: (a) We start with a pseudoline arrangement $L$. (b) We add a triangle containing all intersections of $L$. (c) Each pseudoline is represented by a special edge that is surrounded by a tunnel. (d) Each tunnel consists of tunnel sections. (e) For the crossings of the special edges, we identify parts of the tunnels. (f) We add an apex $u$ and insert triangles to the visible parts of the construction; we enhance the neighborhood of the apex to an essentially 3-connnected graph depicted in (i). Together, (e), (f) and (i) are crucial to enforce the correct combinatorics. (g) In the correctness proof, we use a small sphere around the apex and the projection of each special edge onto the sphere. (h) We will argue that the combinatorics on the sphere are correct and then project the special edges onto a plane. This will give a stretched arrangement. (i) The neighborhood graph of the apex $u$.
On a high level, our construction of $C$ goes along the following lines: We add a helper triangle that contains all intersections of the pseudolines, see Figure 2(b). We place each pseudoline in $\mathbb{R}^{3}$ and replace it by a special edge of the complex $C$. We surround the special edges by tunnels, see Figure 2(c) and (d). For each crossing in $L$, we glue the corresponding tunnel sections together, see Figure 2(e). At last, we insert an apex $u$ high above that is connected to all visible tunnel parts, see Figure 2(f) and we insert additional objects in order to ensure that the neighborhood of $u$ is an essentially 3-connected graph, Figure 2(i).
It is relatively straightforward to verify that if $L$ is stretchable, then the complex $C$ embeds geometrically into $\mathbb{R}^{3}$. The other direction requires more care and work: We show that a geometric embedding of $C$ induces a line arrangement with the same combinatorics as $L$. The idea of the proof is to consider a small sphere around the apex $u$ and to project its neighborhood and the special edges onto the sphere, see Figure 2(g). Because the neighborhood graph of $u$ is essentially 3-connected by construction, all its crossing-free drawings on the sphere are equivalent. This is an crucial property to show that each special edge lies in the projection of its tunnel roof (when restriction the attention to an interesting part within the helper triangle). We remark, that our proof does not show this explicitly. Instead, we establish some even stronger properties. As a consequence, the projection of the tunnels have the intended combinatorics and thus also the special edges. At last, we project the arcs from the sphere onto a plane, see Figure 2(h). In this way, we obtain a line arrangement with the same combinatorics as $L$.
In order to show hardness of $\textsc{GEM}_{3\rightarrow 3}$}, we use a similar construction, in which we “fatten” each triangle to a tetrahedron, by adding extra vertices.
We finally present a dimension reduction, i.e., we reduce $\textsc{GEM}_{k\rightarrow d}$} to $\textsc{GEM}_{k+1\rightarrow d+1}$}. Given a $k$-complex $C$, we create a $(k+1)$-complex $C^{+}$ that contains $C$ and has two additional vertices $a$ and $b$. Moreover, for each subset $e$ of $C$, $C^{+}$ has the additional subsets $e\cup\{a\}$ and $e\cup\{b\}$. We prove that $C$ geometrically embeds in $\mathbb{R}^{d}$ if and only if $C^{+}$ geometrically embeds in $\mathbb{R}^{d+1}$.
In this way, we show that distinguishing PL embeddable and geometrically embeddable complexes is $\exists\mathbb{R}$-complete.
1.1 Existential Theory of the Reals
The class of the existential theory of the reals $\exists\mathbb{R}$ (pronounced as ‘exists R’, ‘ER’, or ‘ETR’) is a complexity class which has gained a lot of interest in recent years, specifically in the computational geometry community. To define this class, we first consider the algorithmic problem Existential Theory of the Reals (ETR). An instance of this problem consists of a sentence of the form
$\exists x_{1},\ldots,x_{n}\in\mathbb{R}:\Phi(x_{1},\ldots,x_{n}),$
where $\Phi$ is a well-formed quantifier-free formula over the alphabet $\{0,1,+,\cdot,\geq,>,\wedge,\vee,\neg\}$, and the goal is to check whether this sentence is true. As an example of an ETR-instance, consider $\exists x,y\in\mathbb{R}\colon\Phi(x,y)=(x\cdot y^{2}+x\geq 0)\wedge\neg(y<2x)$, for which the goal is to determine whether there exist real numbers $x$ and $y$ satisfying the formula $\Phi(x,y)$.
The complexity class $\exists\mathbb{R}$ is the family of all problems that admit a polynomial-time many-one reduction to ETR. It is known that
$\textrm{NP}\subseteq\exists\mathbb{R}\subseteq\textrm{PSPACE}.$
The first inclusion follows from the definition of $\exists\mathbb{R}$. Showing the second inclusion was first established by Canny in his seminal paper [undefp]. The complexity class $\exists\mathbb{R}$ gains its significance because a number of well-studied problems from different areas of theoretical computer science have been shown to be complete for this class.
Famous examples from discrete geometry are the recognition of geometric structures, such as unit disk graphs [undefak], segment intersection graphs [undefah], Stretchability [undefao, undefaaa], and order type realizability [undefah]. Other $\exists\mathbb{R}$-complete problems are related to graph drawing [undefag], Nash-Equilibria [undeff, undefz], geometric packing [undefd], the art gallery problem [undefb], non-negative matrix factorization [undefaz], polytopes [undefu, undefav], geometric linkage constructions [undef], training neural networks [undefc], visibility graphs [undefr], continuous constraint satisfaction problems [undefan], and convex covers [undefa]. The fascination for the complexity class stems not merely from the number of $\exists\mathbb{R}$-complete problems but from the large scope of seemingly unrelated $\exists\mathbb{R}$-complete problems. We refer the reader to the lecture notes by Matoušek [undefah] and surveys by Schaefer [undefaw] and Cardinal [undefq] for more information on the complexity class $\exists\mathbb{R}$.
1.2 Definitions
Simplex.
A $k$-simplex $\sigma$ is a $k$-dimensional polytope which is the convex hull of its $k+1$ vertices $V$, which are not contained in the same (k1)(k-1)-dimensional hyperplane. Hence, a 0-simplex corresponds to a point, a 1-simplex to a segment, and a 2-simplex to a triangle etc. The convex hull of any nonempty proper subset of $V$ is called a face of $\sigma$. A simplicial complex $K$ is a set of simplices satifying the following two conditions: (i) Every face of a simplex from $K$ is also in $K$. (ii) For any two simplices $\sigma_{1},\sigma_{2}\in K$ with a non-empty intersection, the intersection $\sigma_{1}\cap\sigma_{2}$ is a face of both simplices $\sigma_{1}$ and $\sigma_{2}$. The purely combinatorial counterpart to a simplicial complex is an abstract simplicial complex, which we refer to simply as a complex.
Complex.
A complex $C=(V,E)$ is a finite set $V$ together with a collection of subsets $E\subseteq 2^{V}$ which is closed under taking subsets, i.e., $e\in E$ and $e^{\prime}\subseteq e$ imply that $e^{\prime}\in E$. A $k$-complex is a complex where the largest subset contains exactly $k+1$ elements. We call a complex pure if all (inclusion-wise) maximal elements in $E$ have the same cardinality.
For any vertex $v\in V$ in a $k$-complex $C=(V,E)$, the neighbourhood of $v$ gives rise to a lower dimensional complex $C_{v}:=(V^{\prime},E^{\prime})$, where $E^{\prime}:=\{e\setminus\{v\}\mid v\in e\in E\}$ and $V^{\prime}:=N(v)=\bigcup_{e\in E^{\prime}}e$ are the neighbors of $v$. Complexes are in close relation to Hypergraphs.
Hypergraphs.
Hypergraphs generalize graphs by allowing edges to contain any number of vertices. Formally, a hypergraph $H$ is a pair $H=(V,E)$ where $V$ is a set of vertices, and $E$ is a set of non-empty subsets of $V$ called hyperedges (or edges). A $k$-uniform hypergraph is a hypergraph such that all its hyperedges contain exactly $k$ elements. Note that the maximal sets of a pure $k$-complex yield a $(k+1)$-uniform hypergraph and vice versa. Hence, $(k+1)$-uniform hypergraphs and pure $k$-complexes are in a straight-forward one-to-one correspondence. A simplicial representation of a $(k+1)$-uniform hypergraph is a geometric embedding of the corresponding complex.
Geometric embeddings.
A geometric embedding of a complex $C=(V,E)$ in $\mathbb{R}^{d}$ is a function $\varphi\colon V\rightarrow\mathbb{R}^{d}$ fulfilling the following two properties: (i) for every $e\in E$, $\overline{\varphi}(e):={\textrm{conv}}(\{\varphi(v)\colon v\in e\})$ is a simplex of dimension |e|1|e|-1 and (ii) for every pair $e,e^{\prime}\in E$, it holds that
$\overline{\varphi}(e)\cap\overline{\varphi}(e^{\prime})=\overline{\varphi}(e\cap e^{\prime}).$
Note that if $\varphi$ is a geometric embedding, then $\{\overline{\varphi}(e):e\in E\}$ is a simplicial complex. The problem $\textsc{GEM}_{k\rightarrow d}$ asks whether a given $k$-complex has a geometric embedding in $\mathbb{R}^{d}$.
Topological and PL embeddings.
Consider a complex $C=(V,E)$. In contrast to geometric embeddings, for PL or topological embeddings it is not sufficient to describe the mapping of the vertices $V$. Choose $d^{\prime}$ so large that $C$ admits a geometric embedding $\varphi^{\prime}:V\to\mathbb{R}^{d^{\prime}}$, and define $S=\bigcup_{e\in E}\overline{\varphi^{\prime}}(e)$. We then say that an injective and continuous function $\varphi:S\to\mathbb{R}^{d}$ is a topological embedding of $C$ in $\mathbb{R}^{d}$. If furthermore for each $e\in E$, the image $\varphi(\overline{\varphi^{\prime}}(e))$ is a finite union of connected subsets of (|e|1)(|e|-1)-dimensional hyperplanes, then $\varphi$ is a piecewise linear (PL) embedding. The problem $\textsc{Embed}_{k\rightarrow d}$ asks whether a given $k$-complex has a PL embedding in $\mathbb{R}^{d}$.
Graph Drawings.
A graph is a 1-complex. A graph is planar if there exists a crossing-free drawing in the plane, i.e., a (topological) embedding in $\mathbb{R}^{2}$. As mentioned above, a graph has a topological embedding in $\mathbb{R}^{2}$ if and only if it has a geometric embedding in $\mathbb{R}^{2}$. A plane graph is a planar graph together with a specified crossing-free drawing. By means of stereographic projection, any graph that has a crossing-free drawing in the plane also has a crossing-free drawing on the sphere and vice versa. Two drawings of a graph (in the plane or on the sphere) are equivalent if they can be transformed into one another by a homeomorphism (of the plane or the sphere). In particular, two equivalent drawings have the same set of faces. Consider a plane graph $G$ and let $D^{\prime}$ be the specified drawing of $G$. When talking about an (arbitrary) drawing $D$ of $G$, we always mean that $D$ is equivalent to $D^{\prime}$.
Stretchability.
A pseudoline arrangement is a family of curves that apart from ‘straightness’ share similar properties with a line arrangement. More formally, a (Euclidean) pseudoline arrangement is an arrangement of $x$-monotone curves in the Euclidean plane such that any two meet in exactly one point. In fact, each pseudoline arrangement can be encoded by a wiring diagram; see also Figure 4(a). A pseudoline arrangement is stretchable if it is combinatorially equivalent to an arrangement of straight-lines, i.e., if the arrangements can be transformed into one another by a homeomorphism of the plane. Stretchability denotes the algorithmic problem of deciding whether a given pseudoline arrangement is stretchable. In a seminal paper, Shor [undefaaa] proved that Stretchability is complete for the existential theory of the reals; for a stream-line exposition of this result see the expository paper by Matoušek [undefah].
1.3 Pitfalls
While the general proof ideas are fairly straightforward, our arguments in Section 2 may at first glance appear a bit tedious. In the following, we highlight one of the appearing challenges. It is easy to see that each special edge lies inside its tunnel in any geometric embedding. It follows that the projection of the special edge lies also inside the projection of the tunnel on the sphere centered at the apex. Furthermore, we know that the roof of the tunnels are seen by the apex. One may be tempted to (directly) conclude that the projection of the special edge is thus also contained in the projection of the roof; the underlying thought being that the projection of the tunnel bottom lies below the tunnel roof in the geometric representation and thus the projection of the tunnel bottom is contained in the projection of the tunnel roof. Yet, the latter is not true in general, as can be seen in Figure 3. In the figure, the tunnel bottom is not covered by the roof. We (implicitly) show that the projection of the special edge lies inside the projection of the roof by establishing some even stronger topological and geometric properties.
Figure 3: From the perspective of $u$, the tunnel bottom is not always hidden below the tunnel roof. From the three tunnel sections displayed, the bottom (yellow) of the middle section is partially visible from the apex.
2 The Proof
In this section, we prove Theorem 1. Our proof consists of the following three parts.
  1. a)
    Establishing $\exists\mathbb{R}$-membership (Section 2.1: Lemma 3).
  2. b)
    Showing $\exists\mathbb{R}$-hardness in $\mathbb{R}^{3}$, i.e., of $\textsc{GEM}_{2\rightarrow 3}$ and $\textsc{GEM}_{3\rightarrow 3}$ (Section 2.2: Theorems 4 and 5).
  3. c)
    Reducing $\textsc{GEM}_{k\rightarrow d}$ to $\textsc{GEM}_{k+1\rightarrow d+1}$ (Section 2.3: Lemma 6).
Together Lemmas 3, 4, 5 and 6 prove Theorem 1.
2.1 Membership
In this subsection, we show $\exists\mathbb{R}$-membership of $\textsc{GEM}_{k\rightarrow d}$. Note that this is essentially folklore [undefm]. We present a proof for the sake of completeness.
Lemma 3.
For all $k,d\in\mathbb{N}$, the decision problem $\textsc{GEM}_{k\rightarrow d}$ is contained in $\exists\mathbb{R}$.
Proof.
In order to show membership in $\exists\mathbb{R}$, we use the following characterization by Erickson, Hoog and Miltzow [undefv]: A problem $P$ lies in $\exists\mathbb{R}$ if and only if there exists a real verification algorithm $A$ for $P$ that runs in polynomial time on the real RAM. In particular, for every yes-instance $I$ of $P$ there exists a polynomial sized witness $w$ such that $A(I,w)$ returns yes, and for every no-instance $I$ of $P$ and any witness $w$, $A(I,w)$ returns no. In contrast to the definition of the complexity class NP, we also allow witnesses that consist of real numbers. Consequently, we execute $A$ on the real RAM as well.
It remains to present a real verification algorithm for $\textsc{GEM}_{k\rightarrow d}$. While the witness describes the coordinates of the vertices, the algorithm checks for intersections between any two simplices. Note that each simplex is a convex set and the intersection of convex sets is a convex set as well. For any simplex $S$ with $n$ vertices, we can efficiently determine $n$ linear inequalities and at most one linear equality that together describe $S$. Then checking for intersections can be reduced to a linear program, which is polynomial time solvable. This finishes the description of the real verification algorithm. ∎
2.2 Hardness in three dimensions
This section is dedicated to proving Theorem 1 for $d=3$ and $k\in\{2,3\}$. The crucial part lies in the case $k=2$. For the benefit of the reader, we include a glossary in Table 3.
Theorem 4.
The decision problem $\textsc{GEM}_{2\rightarrow 3}$ is $\exists\mathbb{R}$-hard.
Proof.
We reduce from the $\exists\mathbb{R}$-hard problem Stretchability, as described in Section 1.2. Let $L$ be an arrangement of $n$ pseudolines in the plane. Every pseudoline arrangement has a representation as a wiring diagram in which each pseudoline is given by a monotone curve consisting of 2n12n-1 sections. For an illustration consider Figure 4(a); each section could be represented by a segment, however for a visual appealing display, the bend points are rounded. We add a pseudoline $\ell_{0}$ that intersects all pseudolines in the beginning, see Figure 4(a), and call the resulting pseudoline arrangement $L^{*}$. Note that $L^{*}$ is stretchable if and only if $L$ is stretchable. For later reference, we endow a natural orientation upon each pseudoline from left to right.
(a) A pseudoline arrangement $L^{*}$.
(b) The crossing diagram contains an additional helper triangle.
Figure 4: Adding an extra pseudoline $\ell_{0}$ and the helper triangle $\triangle$ to the construction.
In the following, we construct a 2-complex $C=(V,E)$ that allows for a geometric realization if and only if $L^{*}$ (and thus $L$) is stretchable. In order to define $C$, we add a helper triangle $\triangle$ to our arrangement that intersects the pseudolines of $L^{+}$ as illustrated in Figure 4(b).
Construction of the 2-complex.
In order to define $C$, we use the intended geometric embedding. We will refer to the subsets in $C$ as vertices, edges, and triangles depending on whether they contain one, two or three elements. The construction has five steps.
In the first step, we place the pseudolines and the helper triangle $\triangle$ in 3-space. Each pseudoline $\ell_{i}$ lies in the plane $z=i$ such that an observer high above (at infinity) sees the wiring diagram. Similarly, we place the segments of the helper triangle $\triangle$ in 3-space such that it lies in the plane $z=n+1$. Note that no two pseudolines intersect. Therefore, we can surround each lifted pseudoline by a triangulated sphere which we call a tunnel; see also Figure 5(a). The tunnel $T^{+}_{i}$ of $\ell_{i}$ is formed by $2n+3+i$ sections; later, we will be particularly interested in a part of a tunnel, denoted by $T_{i}$, in which the first two and last two sections are removed. Each section consists of six triangles forming a triangulated triangular prism as illustrated in Figure 5(b). We close the tunnel with triangles at the ends and think of the bottom side of the prism to lie in the plane $z=i-\nicefrac{{1}}{{2}}$ (for now). The remaining part of the tunnel, i.e., the tunnel without its bottom, constitutes the roof, see Figure 5(c). The roof contains three disjoint paths of length $2n+4+i$. The edges and vertices on the boundary of the bottom and the roof form the left and right roof path, respectively, when deleting the edges of the closing triangles. The remaining vertices induce the central roof path. The three roof paths are thickened in Figures 5(a) and 5(c).
(a) A tunnel viewed from side.
(b) A tunnel section.
(c) A tunnel viewed from above.
Figure 5: First step in the construction of the complex $C$ – tunnel construction.
Note that we do not add a tunnel for the helper triangle. We distribute the sections along $T^{+}_{i}$ to edges and crossings of the crossing diagram as follows: Generally, we associate one section per edge and one section per crossing of two pseudolines. Moreover, we associate one extra section of $T^{+}_{j}$ to a crossing of $\ell_{i}$ and $\ell_{j}$ whenever $i<j$. In order to represent the pseudoline $\ell_{i}$, we insert a special edge $e_{i}$ between the two top vertices on either end of the tunnel; for later reference, we denote the start vertex by $s_{i}$ and the end vertex by $t_{i}$. The special edge $e_{i}$ is intended to lie inside the tunnel.
In the second step, we identify parts of the tunnels. To this end, consider the tunnel sections assigned to a crossing of a pseudoline $\ell_{i}$ with $\ell_{j}$, $i<j$. Recall that we assigned one section of $T^{+}_{i}$ and two sections of $T^{+}_{j}$ to the crossing. We identify the four triangles in the bottom of the two sections of $T^{+}_{j}$ with the four triangles in the roof of one section of $T^{+}_{i}$ as indicated in Figure 6. Note that we hereby identify six vertices, four of which belong to a left or right roof path of both, $T^{+}_{i}$ and $T^{+}_{j}$.
Figure 6: Second step in the construction of the complex $C$: gluing of tunnel parts.
In the third step, we add a new vertex to the construction that we call the apex and denote it by $u$. We think of $u$ as the observer high above (at infinity) and insert a triangle defined by $u$ and the vertices of every tunnel edge that is visible from $u$. Note that every roof section that is neither glued in a crossings nor hidden by the helper triangle is visible. Moreover, no bottom of any tunnel is visible in the intended geometric embedding.
In the fourth step, we enhance the $1$-complex induced by the neighborhood $N(u)$ of the apex $u$ such that it corresponds to an essentially $3$-connected planar graph $G^{+}$. We call a graph essentially $3$-connected if it is a subdivision of a $3$-connected graph. With the description so far, the $1$-complex corresponds to the graph $H$ depicted in black in Figure 7.
Figure 7: Third and fourth step in the construction of the complex $C$: neighborhood of the apex $u$. The graph $H$ depicted in black after the third step. Together with the gray edges, the graph is a candidate for the essentially 3-connected plane graph $G^{+}$ and it subgraph $G$ inside $\triangle$.
To construct $G^{+}$, we make use of the following fact.
Claim 1.
For every plane graph $G_{1}=(V_{1},E_{1})$, there exists an essentially 3-connected plane graph $G_{2}=(V_{2},E_{2})$ such that $G_{1}$ is a subgraph of $G_{2}$ and any straight-line drawing $D_{1}$ of $G_{1}$ in the plane can be extended to a straight-line drawing of $G_{2}$. Moreover, if the maximum face degree of $G_{1}$ is $k$, then the size of $G_{2}$ can be bounded by $|V_{2}|+|E_{2}|\leq O(k|V_{1}|)$.
Proof.
In order to construct $G_{2}$, we start with a drawing of $G_{1}$. First, we ensure that $G_{1}$ has at least 4 vertices. Second, we guarantee 2-connectedness by inserting edges; the edges are represented by potentially non-straight curves. To this end, we iteratively insert edges between different connected components; this ensures connectedness. Afterwards we insert one vertex in each face and insert one edge to each incident vertex of this face; this ensures 2-connectedness. Third, we triangulate each face by repeating the last step: we insert a vertex into each face and insert an edge to every vertex incident to this face. By 2-connectedness, we obtain a triangulation $T$ on at least four vertices and hence, a 3-connected plane graph. Lastly, we subdivide each new edge so that the number of subdivision vertices equals to the degree of the face in $G_{1}$ in which the edge has been inserted. The resulting graph is $G_{2}$ and by construction essentially 3-connected. Note that $T$ has $O(|V_{1}|)$ vertices. Therefore, $G_{2}$ has $O(k|V_{1}|)$ edges.
It remains to show that any straight-line drawing $D_{1}$ of the plane graph $G_{1}$ can be extended to a straight-line drawing of the planar graph $G_{2}$. Because $D_{1}$ and $G_{1}$ have the same set of faces, we can insert the additional edges of $T$ by polylines. Following the face boundary, the number of bends on each edge is upper bounded by the degree of the face of $G_{1}$ in which it is inserted. Hence, we can easily extend $D_{1}$ to a straight-line drawing of $G_{2}$. ∎
Let $G^{+}:=G_{2}$ be an essentially 3-connected plane graph guaranteed by Claim 1 for the case that $G_{1}=H$. Note that $G_{1}$ has $O(n^{2})$ vertices and edges, and every face has degree $O(n)$. Hence, the size of $G_{2}$ is in $O(n^{3})$. We denote the outer face of $G_{2}$ by $f_{0}$. The reader is invited to think about the far more sparse graph depicted in Figure 7, which also serves as a candidate for $G^{+}$. Indeed, the depicted graph also fulfills all properties necessary for our construction; however, not all properties of Claim 1. For example, the depicted graph is even $3$-connected. The proof of this is straightforward, but a bit tedious. Thus, we leave it as an exercise to the interested reader to check that the graph remains connected even after the deletion of any two vertices or alternatively, that any pair of vertices is connected by three disjoint paths.
Later, the subgraph $G$ of $G^{+}$ that is induced by all vertices of $\bigcup_{i}T_{i}$ will be of particular interest; in Figure 7, these vertices (and their convex hull) lie inside the helper triangle $\triangle$. Recall that $T_{i}$ denote the part of the tunnel $T^{+}_{i}$ obtained by deleting the first two and last two sections.
It is a well-known fact that all (straight-line or topological) planar drawings of a 3-connected planar graph on the sphere are equivalent [undefae]; for a definition of equivalent drawings consult Section 1.2. Consequently, the result extends to essentially $3$-connected graphs as it also holds for topological drawings. For later reference, we note the following.
Claim 2.
The planar graph $G^{+}$ is essentially $3$-connected. Therefore, all crossing-free drawings of $G^{+}$ on a sphere are equivalent. Furthermore, any straight-line drawing of $H$ in the plane can be extended to a straight-line drawing of $G^{+}$.
We ensure that the neighborhood complex of $u$ is the underlying planar graph of $G^{+}$, i.e, for each edge of $G^{+}$ not present in $H$, we insert a triangle formed by the vertices of this edge together with $u$ and call the resulting complex $\overline{C}$.
In the fifth and last step, our final complex $C$ consist of two copies of $\overline{C}$ in which the apex vertices are identified. We will later use these two copies in order to guarantee that in any geometric embedding the apex lies outside of all tunnels for one copy of $\overline{C}$. This finishes the construction of the complex $C$.
Time Complexity.
In order to verify that the construction shows $\exists\mathbb{R}$-hardness, we argue that it has a running time that is polynomial in the size of the input. To this end, note that a pseudoline arrangement with $n$ pseudolines can be described by the order of the $O(n^{2})$ crossings. Thus, the input size is $N=O(n^{2})$. After adding the helper triangle and $\ell_{0}$, the crossing diagram still has a size in $O(n^{2})$. It is easy to see that our construction has a size proportional to $N^{3/2}$: For each segment and crossing of the diagram, we build a constant size construction involving the apex. Moreover, we add a triangle for every (additional) edge in $G^{+}$; recall that $G^{+}$ has size $O(n^{3})$. Consequently, the total construction has size $O(n^{3})=O(N^{3/2})$. We remark, that a more careful choice of $G^{+}$, as in Figure 7, yields a construction that is linear in $N$.
Correctness.
It remains to show that the pseudoline arrangement $L$ is stretchable if and only if $C$ has a geometric embedding in $\mathbb{R}^{3}$.
Correctness I: Stretchability implies Embedabbility.
If $L$ is stretchable, it is relatively straight-foward to construct a geometric embedding of $C$.
Claim 3.
If $L$ is stretchable, then $C$ has a geometric embedding.
Proof.
The construction of the geometric embedding goes along the same lines as the construction of $C$. Consider the stretched line arrangement $L^{\prime}$ equivalent to the pseudoline arrangement $L$, we construct $L^{\prime\prime}$, by adding $\ell_{0}$ to the left of all crossings of $L^{\prime}$.
Next we show how to add the helper triangle $\triangle$ such that the intersection pattern is as depicted in the crossing diagram. Note that we merely have to construct the triangle in a way that all intersections of $L^{\prime\prime}$ are contained in it and the vertices of the helper triangle lie in the correct faces. For an illustration of the following argument see Figure 8.
Figure 8: Illustration for the proof of Claim 2 concerning the insertion of $\triangle$. By scaling $L^{\prime\prime}$ while keeping point $p$ and the helper triangle fixed, we can ensure that all intersections lie inside the helper triangle eventually.
The helper triangle has three vertices $v_{0},v_{1},v_{2}$. By construction, the two vertices $v_{0},v_{1}$ are supposed to lie in the bottom cell of $L^{*}$ which is bounded by $\ell_{0}$ and $\ell_{n}$ and potentially other pseudolines. Similarly, $v_{2}$ is in the upper cell bounded exactly by $\ell_{0}$ and $\ell_{n}$. Let us fix some triangle $\triangle$ that contains the intersection point $p$ of $\ell_{0}$ and $\ell_{n}$ with the properties above. Consider the (gray) region that contains all intersections of $L^{\prime\prime}$. By scaling $L^{\prime\prime}$ while fixing $p$ and the triangle $\triangle$, the gray region becomes arbitrarily small and is eventually contained in $\triangle$. This shows that the helper triangle can be added as wished.
Afterwards, we lift the lines and the helper triangle in $3$-space, construct the tunnels, and enhance $H$ to $G^{+}$. For the latter step, we use the fact that we can extend any straight-line drawing of $H$ to a straight-line drawing of $G^{+}$ by Claim 2. (In case that we want to use $G^{+}$ as in Figure 7, we need to use the fact that the drawing of $H$ comes from the stretched line arrangement and thus the faces are convex.) Afterwards, we insert the apex $u$ (high enough) above, make a second copy by taking the mirror image, and identify the apices. This yields a geometric embedding of $C$. ∎
Correctness II: Embedabbility implies Stretchability.
The reverse direction is more involved. Let $\varphi$ denote a geometric embedding of $C$. To show that $L$ is stretchable, we start with a collection of crucial properties.
By definition, each tunnel forms a closed topological sphere; all of which are pairwise disjoint. By a generalization of the Jordan curve theorem, also known as the Jordan–Brouwer separation theorem, any topological embedding of a (d1)(d-1)-sphere in $\mathbb{R}^{d}$ splits the space into two components [undefaal]; we refer to the bounded component of a tunnel or any other topological sphere as its inside and to the unbounded component as its outside.
Claim 4.
There exists a copy $\overline{C}$ in $C$ such that the apex $u$ lies outside all tunnels of $\varphi(\overline{C})$.
Proof.
Note that the apex $u$ lies inside at most one tunnel of $C$; otherwise, among all tunnels containing $u$, the innermost separates $u$ from the outermost. A contradiction to the fact that shares and edge with the vertices of all tunnels. Consequently, for at least one copy of $\overline{C}$ contained in $C$, the apex $u$ lies outside of each tunnel. ∎
From now on, we focus on the geometric embedding $\varphi(\overline{C})$ of this $\overline{C}$ and do not make further use of the other copy.
Claim 5.
In $\varphi(\overline{C})$, the special edge $e_{i}$ lies inside its tunnel $T^{+}_{i}$ for all $i$.
Proof.
Consider the vertex $s_{i}$ of $e_{i}$ and the tunnel $T^{+}_{i}$ forming a closed topological sphere incident to $s_{i}$. By construction, all roof edges of the first section of $T^{+}_{i}$ are visible by the apex $u$ and thus form triangles with $u$. Note that the edges non-incident to $s_{i}$ contain a cycle, see Figure 9(a).
(a) A top view of first section.
(b) The tent over $s_{i}$.
Figure 9: Illustration for the proof of Claim 5: the spheres surrounding $s_{i}$.
The triangles between the apex $u$ and this cycle form a tent on top of $T^{+}_{i}$, see Figure 9(b). In particular, the tent together with the tunnel roof of the first section form another sphere incident to $s_{i}$. By Claim 4, $u$ is outside the tunnel. If $e_{i}$ started towards the outside of the tunnel, then it would be trapped inside the tent-sphere which is not incident to $t_{i}$. Consequently, $e_{i}$ lies inside the tunnel. ∎
Consider a special edge $e_{i}$ and the wedge $W_{i}$ defined by rays originating at $u$ through points in $e_{i}$, see Figure 10. We say a point $p$ of $(W_{i}\setminus e_{i})$ lies above $e_{i}$ if the segment $up$ does not intersect $e_{i}$; we write $p>e_{i}$. Similarly, $p$ lies below $e$ if the segment $up$ does intersect $e_{i}$; we write $p<e_{i}$.
Figure 10: The wedge $W_{i}$ of $e_{i}$ containing a point $p$ above $e_{i}$ and a point $q$ below $e_{i}$.
Claim 6.
Consider a special edge $e_{i}$. For every two points $p,q$ on $W_{i}\setminus e_{i}$ that belong to a common triangle in $T^{+}_{i}$, the above/below-relation is consistent, i.e., $p<e_{i}\iff q<e_{i}$.
Proof.
Suppose for a contradiction that $p>e_{i}$ and $q<e_{i}$. By visibility of $s_{i}$ and $t_{i}$, $p$ lies neither on the segment $us_{i}$ nor on the segment $ut_{i}$. Hence, the segment $pq$ intersects $e_{i}$ in an inner point, i.e., it intersects $e-\{s_{i},t_{i}\}$, see Figure 10. However, no triangle of $C$ contains $e_{i}$ and hence $e_{i}-\{s_{i},t_{i}\}$ and $pq$ must have an empty intersection. A contradiction. ∎
Now, we consider a small sphere $S$ around the apex $u$ which has no other vertex inside, see Figure 11(a). For each triangle in $\varphi(\overline{C})$ containing $u$, the intersection with the sphere $S$ yields an arc of a great circle. Consequently, we obtain a crossing-free drawing $D^{+}$ of the essentially 3-connected planar graph $G^{+}$ (the neighborhood complex of $u$) with arcs of great circles on $S$. Moreover, for each special edge $e_{i}$, we consider the projection $a_{i}$ of the (artificial) triangle $\{s_{i},t_{i},u\}$ onto the sphere $S$; note that $\{s_{i},t_{i},u\}$ is not a triangle of the complex, see Figure 11(b).
Figure 11: Obtaining a line arrangement from a geometric embedding.
This yields a set of arcs $\mathcal{A}$. In the remainder, we show that $\mathcal{A}$ has the same combinatorics/intersection pattern as the pseudoline arrangement $L^{*}$, i.e., the order of intersections along each arc/pseudoline are the same. Afterwards, we project $\mathcal{A}$ to a plane as illustrated in Figure 11(c) and obtain a line arrangement proving that $L^{*}$ (and thus $L$) is stretchable.
Let $D$ denote the restriction of the drawing $D^{+}$ to the graph $G$. By Claim 2, all crossing-free drawings of $G^{+}$ on the sphere are equivalent. This implies the following fact:
Claim 7.
In $D^{+}$, the helper triangle separates the subdrawing $D$ from the vertices $\{s_{j},t_{j}\mid j=0,\dots,n\}$. In particular, by convexity, the convex hull of any vertex subset $U$ in $D$ is contained in the helper triangle.
For a tunnel, we call the intersection of two consecutive sections a tunnel loop. Note that each tunnel loop consists of three edges, two of which belong to the roof of the tunnel and form a rafter; the remaining edge belongs to the bottom of the tunnel.
Claim 8.
$D\cup\mathcal{A}$ fulfills the following properties for all $i$:
  1. i)
    The arc $a_{i}$ intersects each rafter of $T_{i}$ exactly one time.
  2. ii)
    The intersection points appear in the correct order along $a_{i}$, i.e., they follow the natural order along the tunnel.
  3. iii)
    The faces of a roof section in $T_{i}$ cover an interval of $a_{i}$.
Proof.
Consider a tunnel loop o{o} of tunnel $T^{+}_{i}$ and its (projected) vertices in $D$. We first show that the arc $a_{i}$ intersects at least one (potentially artificial) segment connecting the vertices of o{o} in $D$. To this end, we may assume that $o$ is disjoint from $s_{i}$ and $t_{i}$ and consider the restriction of $\varphi(\overline{C})$ to the plane $P_{i}$ containing $u$ and $e_{i}$. By Claims 4 and 5, the apex $u$ is outside of the tunnel $T^{+}_{i}$ while $e_{i}$ is inside $T^{+}_{i}$. Hence, there exist curves $\gamma_{a},\gamma_{b}\subset T^{+}_{i}$ from $s_{i}$ to $t_{i}$ on $P_{i}$ that together enclose $e_{i}$ and separate $e_{i}$ from $u$, see Figure 12.
Figure 12: An illustration for the proof of Claim 8: the curve $\gamma_{a}$ in $P_{i}$ separates $e_{i}$ from $u$.
Because $s_{i}$ and $t_{i}$ are visible from $u$, the curves intersect the segments $s_{i}u$ and $t_{i}u$ exactly in $s_{i}$ and $t_{i}$, respectively. We denote the curve contained in the triangle of $u,s_{i},t_{i}$ by $\gamma_{a}$. The curve $\gamma_{a}$ contains a point $x\in{o}$ because $s_{i}$ and $t_{i}$ lie in different components of $T^{+}_{i}-{o}$. Hence, the ray supporting $ux$ intersects $e_{i}$, i.e., $x>e_{i}$.
Next we show that if o{o} belong to $T_{i}$, then the curve $\gamma_{b}$ intersects o{o} in a point below $e_{i}$. Analogous to $\gamma_{a}$, the curve $\gamma_{b}$ also contains a point $x^{\prime}\in{o}$. It remains to show that $x^{\prime}$ is below $e_{i}$. Note that this does not hold, only in the case that $x^{\prime}$ is to the left of the ray $us_{i}$ or to the right of the ray $ut_{i}$; otherwise, $\gamma_{b}$ is below $e_{i}$. By Claim 7, the projection of the helper triangle separates $D$ from the vertices $s_{i},t_{i}$ in $D^{+}$. Consequently, the helper triangle $\triangle$ and $P_{i}$ intersect in two points $q_{1}$ and $q_{2}$. The points $q_{1}$ and $q_{2}$ must be above $e_{i}$, because $u$ forms a triangle with every edge of $\triangle$. By Claim 7 and the fact that the vertices of $o$ are contained in $D$, the vertices of $o$ and their convex hull are inside $\triangle$ in $D$, see Figure 13. Note that $o$ is not completely contained in $D$ because it contains an edge of the tunnel bottom.
Figure 13: An illustration for the proof of Claim 8. The wedge $W_{i}$ and the helper triangle $\triangle$ (yellow) intersect in the segment $q_{1}q_{2}$. We drew the projection of $D$ onto the the helper triangle. In $D$, the vertices of the sloop $o$ (purple) are contained in $\triangle$. Therefore, the point $x^{\prime}$ lies between the rays $uq_{1}$ and $uq_{2}$ on $W_{i}$ and thus below $e_{i}$.
This implies that $x^{\prime}$ is in a part of $\gamma_{b}$ that is bounded by the rays $uq_{1}$ and $uq_{2}$, see Figure 12. Therefore, $x^{\prime}$ lies below $e_{i}$ and is thus invisible from $u$. It follows that $x^{\prime}$ belongs to an invisible edge of o{o}, i.e., $x^{\prime}$ is contained in the edge of $o$ belonging to the tunnel bottom; we denote it by $b_{o}$. By Claim 6, no point of $b_{o}$ lies above $e_{i}$. Consequently, $x>e_{i}$ belongs to the tunnel roof, i.e., the rafter of o{o}.
Moreover, no triangle of $T_{i}$ incident to $b_{o}$ lies above $e_{i}$. Note that for each triangle in the bottom of $T_{i}$, there exists a tunnel loop o{o} such that $b_{o}$ is incident to it. Hence, no point of the bottom of $T_{i}$ lies above $e_{i}$, i.e., we obtain the following Property 1: the curve $\gamma_{a}$ does not contain points of the bottom of the tunnel $T_{i}$.
It remains to argue that there exists no further intersection point if ${o}\subset T_{i}$. We show that $(\gamma_{a}\cup\gamma_{b})\cap{o}=\{x,x^{\prime}\}$. Suppose that there exists a further point $x^{\prime\prime}\in(\gamma_{a}\cup\gamma_{b})\cap{o}$. Then o{o} must be contained in $P_{i}$ and we may assume that $x,x^{\prime\prime}$ are vertices of o{o}. However, then the edge $xx^{\prime\prime}$ intersects the closed curve formed by concatenating the segments $us_{i}$, $s_{i}t_{i}=e_{i}$, $t_{i}u$ – a contradiction to the properties of a geometric embedding. Consequently, $(\gamma_{a}\cup\gamma_{b})\cap{o}=\{x,x^{\prime}\}$. In particular, we obtain Property 2: $\gamma_{a}$ intersects each rafter of $T_{i}$ exactly once. This implies that $a_{i}$ intersects each rafter in $D$ exactly once and thus proves i).
Together, Properties 1 and 2 imply that $\gamma_{a}$ visits any section of $T_{i}$ exactly once. Consequently, the part of $a_{i}$ covered by the roof faces of any section forms an interval on $a_{i}$. This proves iii).
It remains to prove ii). By Property 2, $\gamma_{a}$ intersects each rafter exactly once. We denote the intersection points on the curve $\gamma_{a}$ with the rafters by $p^{\prime}_{1},\dots,p^{\prime}_{k}$ in the order they appear on $\gamma_{a}$, i.e., $p^{\prime}_{j}$ denotes the intersection of $\gamma_{a}$ with the $j$-th rafter of $T_{i}$. Let $p_{j}$ denote the projection of $p^{\prime}_{j}$ onto $S$, i.e., the intersection point of the $j$-th rafter with $a_{i}$ in $D\cup\mathcal{A}$. We show that $p_{1},\ldots,p_{k}$ appear in this order along $a_{i}$ in $D\cup\{a_{i}\}$, i.e., $p_{j}$ lies before $p_{j+1}$ for all $j$. To this end, we restrict our attention to the wedge $W_{i}\subset P_{i}$ bounded by the rays $uq_{1}$ and $uq_{2}$. For simplicity of the presentation, we transform $W_{i}$ such that (the interesting part of) $e_{i}$ is on the $x$-axis, $s_{i}$ left of $t_{i}$, and that the apex $u$ at infinity above, see also Figure 14(a).
(a) The curve $\gamma_{a}$ shields $e_{i}$ from $u$.
(b) $p^{\prime}_{j+1}$ lies below $\gamma^{\prime}$.
(c) $p^{\prime}_{j+1}$ lies above $\gamma^{\prime}$.
Figure 14: Illustration for the proof of Claim 8 ii).
Suppose for the purpose of a contradiction that $p_{j+1}$ lies left of $p_{j}$, see also Figures 14(b) and 14(c). Note that the initial subcurve of $\gamma_{a}$ until $p^{\prime}_{j}$, denoted by $\gamma_{a}^{\prime}$, shields the initial part of $e_{i}$ until $p_{j}$. If $p^{\prime}_{j+1}$ lies below $\gamma_{a}^{\prime}$ as depicted in Figure 14(b), then $p^{\prime}_{j+1}$ is not visible from $u$. However, by construction, all rafters and thus all points $p^{\prime}_{1},\dots,p^{\prime}_{k}$ are visible from the apex $u$. A contradiction. If $p^{\prime}_{j+1}$ lies above $\gamma_{a}^{\prime}$ as depicted in Figure 14(c), then $p^{\prime}_{i+1}$ is not visible from $u$ because of the subcurve $\gamma_{a}-\gamma_{a}^{\prime}$, a contradiction. Consequently, $p_{j}$ is left of $p_{j+1}$ for all $j$. ∎
In order to define a notion of left and right of $a_{i}$ in $D\cup\{a_{i}\}$, we enhance $D\cup\{a_{i}\}$ by the projection of $\triangle$. The arc $a_{i}$ partitions the interior of (the projection of) $\triangle$ into to two regions. Considering the orientation of $a_{i}$, it is easy to determine if a point within one of these two regions lies left or right.
Every drawing on the sphere can be transformed to a drawing in the plane where any face can be chosen as the outer face. In order to relate notions such as clockwise and counter-clockwise with our drawing on the sphere, we fix the face $f_{0}$ as the outer face of $D^{+}$ such that $v_{0},v_{1},v_{2}$ is a counter clockwise cycle. For $D$, these notions are then inherited because we consider a drawing $D$ that is an induced subdrawing of $D^{+}$.
Claim 9.
In $D\cup\{a_{i}\}$, the vertices of the left and right roof paths of $T_{i}$ are left and right of $a_{i}$.
Proof.
Let $u_{1},\dots,u_{k}$ denote the vertices of the left roof path, $v_{1},\dots,v_{k}$ the vertices of the central roof path, and $w_{1},\dots,w_{k}$ the vertices of the right roof path of $T_{i}$ in $D\cup\{a_{i}\}$. The value of $k$ depends on the number of tunnel sections of $T_{i}$, namely, k=2n4+ik=2n-4+i. For each $j=1,\ldots,k$, it holds that $u_{j},v_{j},w_{j}$ is a rafter which $a_{i}$ intersects exactly once by Claim 8 i). Consequently, if $u_{j}$ lies left of $a_{i}$ then $w_{j}$ lies right of $a_{i}$ and vice versa.
We consider a visible section of $T_{i}$. To this end, let $u_{j}u_{j+1}$ be a visible edge. We show that both $u_{j}$ and $u_{j+1}$ lie left of $a_{i}$ implying that $w_{j}$ and $w_{j+1}$ lie right of $a_{i}$. Because each vertex of the left or right roof path belongs to some visible section, this property implies the claim.
First consider the case that $u_{j}$ and $u_{j+1}$ lie right of $a_{i}$. By Claim 8 ii), the intersection $x_{j}$ of the rafter $u_{j},v_{j},w_{j}$ with $a_{i}$ lies before the intersection $x_{j+1}$ of the rafter $u_{j+1},v_{j+1},w_{j+1}$ with $a_{i}$. However, the cycle $u_{j},u_{j+1},v_{j+1},w_{j+1},w_{j},v_{j}$ is flipped, i.e., it has the outer face $f_{0}$ to its right, however it is supposed to lie to its left, see Figures 15(a) and 15(b). A contradiction.
Figure 15: Illustration for the proof of Claim 9.
Hence it remains to consider the case in which $u_{j}$ and $u_{j+1}$ lie on different sides. We consider the case that $u_{j}$ lies left and $u_{j+1}$ right of $a_{i}$ as illustrated in Figure 15(c). By crossing-freeness, exactly one of the edges $u_{j}u_{j+1}$ and $w_{j}w_{j+1}$ crosses $a_{i}$ between $x_{j}$ and $x_{j+1}$, while the other edge crosses before $x_{j}$ or after $x_{j+1}$ as depicted in Figure 15(c). However, then the roof faces of this section do not cover a consecutive part of $a_{i}$. A contradiction to Claim 8 iii). ∎
Claim 10.
In $\mathcal{A}$, for all $i,j$ with $i<j$, the arc $a_{j}$ crosses the arc $a_{i}$ from left to right.
Proof.
We restrict our attention to the vertices of the left and right roof paths of $T_{i}$ and $T_{j}$. Let $\alpha_{1},\alpha_{2},\alpha_{3}$ and $\beta_{1},\beta_{2},\beta_{3}$ denote the vertices of two rafters of $T_{i}$ such that they are contained in the right and left roof path of $T_{j}$, respectively. For an illustration, consider Figure 16(a). We call their induced rafters, the $\alpha$-rafter and the $\beta$-rafter. By Claim 9, the vertices of a left/right roof path lie on the left/right of its corresponding arc.
(a) The vertices $\alpha_{1},\alpha_{2},\alpha_{3},\beta_{1},\beta_{2},\beta_{3}$.
(b) The arcs $a_{i}$ and $a_{j}$ cross.
(c) Arc $a_{j}$ cannot cross $a_{i}$ from right to left.
Figure 16: Illustration for the proof of Claim 10.
Suppose the arcs $a_{i}$ and $a_{j}$ do not cross. By symmetry, we may assume that $a_{j}$ lies left of $a_{i}$, see Figure 16(b). Recall that the vertices $\alpha_{3}$ and $\beta_{3}$ lie to the right side of $a_{i}$, but on different sides of $a_{j}$. A contradiction.
Now, suppose $a_{j}$ crosses $a_{i}$ from right to left as in Figure 16(c). Then each vertex among $\alpha_{1},\alpha_{3},\beta_{1},\beta_{3}$ lies in one of the four regions defined by $a_{i}$ and $a_{j}$. Moreover $a_{j}$ separates the $\alpha$-rafter of $T_{i}$ containing $\alpha_{1},\alpha_{2}$, and $\alpha_{3}$ from the $\beta$-rafter containing $\beta_{1}$$\beta_{2}$, and $\beta_{3}$. This implies that $a_{i}$ intersects the $\beta$-rafter before the $\alpha$-rafter. A contradiction to Claim 8 ii). ∎
Claim 11.
In $\mathcal{A}$, the order of intersections on each arc $a_{i}$ is the same as for $\ell_{i}$ in $L^{*}$.
Proof.
We consider any three arcs $a_{i},a_{j},a_{k}$ with $i<j<k$. By Claim 10, the vertices $s_{i}$$s_{j}$$s_{k}$$t_{i}$$t_{j}$, and $t_{k}$ lie on the correct sides for all three arcs. It is therefore the case that the order of intersections is correct on one arc of a triple if and only if it is correct on all three arcs. Suppose for a contradiction that order of intersections on $a_{i}$ is not correct, i.e., the intersection of $a_{j}$ and $a_{k}$ lies on the wrong side of $a_{i}$. By symmetry, we consider the case that the intersection is left of $a_{i}$ while it is supposed to lie right of $a_{i}$, see Figure 17.
(a) The subgraph of $G$ induced by vertices of the three tunnels $T_{i},T_{j},T_{k}$.
(b) A wrong order of intersections implies that $a_{i}$ intersects the $\beta$-rafter before the $\alpha$-rafter.
Figure 17: Illustration for the proof of Claim 11.
Let $\alpha_{1},\alpha_{2},\alpha_{3}$ and $\beta_{1},\beta_{2},\beta_{3}$ denote the vertices of two rafters of $T_{i}$ such that they are contained in the left roof path of $T_{j}$ and in the right roof path of $T_{k}$, respectively. For an illustration, consider Figure 17(a). Consequently, by Claim 9, $\alpha_{1},\alpha_{2},\alpha_{3}$ lie left of $a_{j}$ and $\beta_{1},\beta_{2},\beta_{3}$ lie right of $a_{k}$. This implies that the intersection of the $\beta$-rafter with $a_{i}$ lies before the intersection of the $\alpha$-rafter with $a_{i}$. A contradiction to Claim 8 ii). ∎
Claim 12.
In $\mathcal{A}$, all intersections lie in one hemisphere of $S$ bounded by $a_{0}$.
Proof.
Consider the plane $P$ through the arc $a_{0}$ and apex $u$ and assume without loss of generality that $P$ is horizontal. By construction and Claim 11, the first crossing on each arc $a_{i}$ is with the arc $a_{0}$; all other crossings come afterwards. Moreover, Claim 10 ensures that each arc $a_{i}$ crosses $a_{0}$ from left to right. Consequently, all arcs start in one hemisphere of $S$ cut by $P$; let us suppose it is the top hemisphere. Then, the intersection of any two arcs $a_{i}$ and $a_{j}$ lie on the bottom hemisphere; otherwise the arcs contain both intersections of the supporting great circles. ∎
Let $P^{\prime}$ denote a plane obtained by shifting the plane $P$ through $a_{0}$ and $u$ to the bottom, see also Figure 11(c). Then, we ’project’ the arc arrangement onto the plane $P^{\prime}$ by considering the intersection of $P^{\prime}$ with the plane supporting the arc $a_{i}$. Clearly, this intersection yields a line. Hence, we obtain a line arrangement in the plane $P^{\prime}$ that inherits the combinatorics of $\mathcal{A}$ and thus, proves the stretchability of $L$. This finishes the proof of Theorem 4. ∎
Fattening the Complex.
The constructed 2-complex in the proof of Theorem 4 was not pure. Specifically, the special edges are not contained in any triangle. We can obtain a pure 2-complex by adding one new vertex to each special edge such that it forms a special triangle. Similarly, we can add a private vertex to each triangle to form a pure 3-complex $C^{\prime}$. Given a geometric embedding of $C$, these new vertices can easily be added close enough to their defining set in $C$. Hence, $C$ has a geometric embedding if and only if $C^{\prime}$ has a geometric embedding in $\mathbb{R}^{3}$. Together with Theorem 4, this very small modification shows hardness of $\textsc{GEM}_{3\rightarrow 3}$.
Lemma 5.
The decision problem $\textsc{GEM}_{3\rightarrow 3}$ is $\exists\mathbb{R}$-hard.
2.3 Dimension Reduction
In order to show hardness for all remaining cases of Theorem 1, we establish the following dimension reduction. For dimension reductions in the context of PL embeddings, we refer to [undefat, undefaab, undefw, undefas].
Lemma 6.
The decision problem $\textsc{GEM}_{k\rightarrow d}$ reduces to $\textsc{GEM}_{k+1\rightarrow d+1}$.
The idea is to add two apices to a $k$-complex $C$ in order to obtain a $(k+1)$-complex $C^{+}$. We will then argue that $C$ has a geometric embedding in $\mathbb{R}^{d}$ if and only if $C^{+}$ has a geometric embedding in $\mathbb{R}^{d+1}$. More formally, for a complex $C=(V,E)$ and a disjoint vertex set $U$, CUC*U denotes the join complex $(V\cup U,E^{\prime})$ where $E^{\prime}:=\{e\cup u\mid e\in E,u\in U\}$. The following claim immediately implies Lemma 6.
Claim 13.
Let $C=(V,E)$ be a complex, $a,b\notin V$ two new vertices, and $C^{+}:=C*\{a,b\}$ their join complex. Then $C$ has a geometric embedding in $\mathbb{R}^{d}$ if and only if $C^{+}$ has a geometric embedding in $\mathbb{R}^{d+1}$.
Proof.
Let $\varphi$ be a geometric embedding of $C$ in $\mathbb{R}^{d}$. Then, we define for $v\in V\cup\{a,b\}$,
$\varphi^{\prime}(v)=\begin{cases}(\ \ \varphi(v)\ \ ,\ 0\ )&\text{ if }v\in V,\\ (0,\dots,0,+1)&\text{ if }v=a,\\ (0,\dots,0,-1)&\text{ if }v=b.\end{cases}$
It is easy to check that $\varphi^{\prime}$ is a geometric embedding of $C^{+}$ in $\mathbb{R}^{d+1}$: While $a$ and $b$ are well separated in the last coordinate, all other potential intersections happen in the $d$-dimensional subspace induced by the first $d$ coordinates. Hence $\varphi$ implies the correctness of the geometric embedding.
For the reverse direction, consider a geometric embedding $\varphi$ of $C^{+}$ in $\mathbb{R}^{d+1}$. Let $\varphi_{a}:=\varphi(a)$ and $\varphi_{b}:=\varphi(b)$. Without loss of generality, we assume that $\varphi_{a}-\varphi_{b}$ is orthogonal to the first $d$ coordinates, i.e., $\varphi_{a}-\varphi_{b}$ is parallel to the $(d+1)$-st coordinate axis. Let $\overline{\varphi}(C):=\bigcup_{e\in E}\overline{\varphi}(e)$ denote the induced geometric subrepresenation of $C$. We claim that the function $f:\overline{\varphi}(C)\to\mathbb{R}^{d}$ obtained by restricting to the first $d$ coordinates is injective. Thus $\varphi^{\prime}:=f\circ\varphi$ yields a representation of $C$ in $\mathbb{R}^{d}$.
For the purpose of a contradiction, suppose that $f$ is not injective. Then there exist two distinct points $p=(p_{1},\ldots,p_{d+1})$ and $q=(q_{1},\ldots,q_{d+1})$ with $p,q\in\overline{\varphi}(C)$ such that $(p_{1},\ldots,p_{d})=(q_{1},\ldots,q_{d})$ and $p_{d+1}\neq q_{d+1}$. Without loss of generality, we may assume that $p_{d+1}>q_{d+1}$. Consider the plane $P$ spanned by $\varphi_{a},\varphi_{b},p$. Note that $q\in P$. For an illustration, see Figure 18(a).
Figure 18: Illustration for the proof of Claim 13. The geometric embedding $\varphi$ of $C^{+}$, gives a monotone embedding of $C$, otherwise we can find an intersection.
Let us denote with $e_{p}\in E$ and $e_{q}\in E$ any choice of hyperedges such that $p\in\overline{\varphi}(e_{p})$ and $q\in\overline{\varphi}(e_{q})$. Consider the two open segments ${\textrm{seg}}^{\circ}(\varphi_{a},q)\in\overline{\varphi}(e_{q}\cup a)$ and ${\textrm{seg}}^{\circ}(\varphi_{b},p)\in\overline{\varphi}(e_{p}\cup b)$. Clearly, these open segments intersect in a point $x$, as illustrated Figure 18(a). Because $\varphi$ is a geometric embedding, it holds that
$x\in\overline{\varphi}(e_{q}\cup a)\cap\overline{\varphi}(e_{p}\cup b)=\overline{\varphi}(e_{q}\cap e_{p})=\overline{\varphi}(e_{q})\cap\overline{\varphi}(e_{p}).$
[…]
3 Conclusion
We established the computational complexity of $\textsc{GEM}_{k\rightarrow d}$ for all $d\geq 3$ and $k\in\{d-1,d\}$. In particular, we showed that for these values it is complete for $\exists\mathbb{R}$ to distinguish PL embeddable $k$-complexes in $\mathbb{R}^{d}$ from geometrically embeddable ones. Arguably, $\textsc{GEM}_{2\rightarrow 3}$ is the most interesting case.
Investigating the computational complexity for the remaining open entries in Table 2 remains for future work. We strengthen the conjecture of Skopenkov [undefaad] as follows.
Conjecture.
The problem $\textsc{GEM}_{k\rightarrow d}$ is $\exists\mathbb{R}$-complete for all $k,d$ such that $\max\{3,k\}\leq d\leq 2k$.
Acknowledgements.
We thank Arkadiy Skopenkov for his kind and swift help with acquiring literature. We thank Martin Tancer for pointing out a mistake in a previous version of this manuscript. Mikkel Abrahamsen is part of Basic Algorithms Research Copenhagen (BARC), generously supported by the VILLUM Foundation grant 16582. Linda Kleist is generously supported by a postdoc fellowship of the German Academic Exchange Service (DAAD). Tillmann Miltzow is generously supported by the Netherlands Organisation for Scientific Research (NWO) under project no. 016.Veni.192.250.
Table 3: A glossary for notions used in the proof of Theorem 4.
symbol meaning
$L=\{\ell_{1},\ldots,\ell_{n}\}$ pseudoline arrangement
$n$ number of pseudolines in $L$, Figure 4(a)
$\ell_{0}$ additional pseudoline
$^{*}:=L\cup\{\ell_{0}\} LL$ together with $\ell_{0}$
$\triangle$ the helper triangle, all intersections of $L^{*}$ are contained inside
$C=(V,E)$ the simplicial complex that we construct
$T^{+}_{i}$ tunnel around pseudoline $\ell_{i}$, Figure 5(a)
$T_{i}$ tunnel $T^{+}_{i}$ without first and last two sections
section part of tunnel, Figure 5(b)
tunnel bottom the part of tunnel $T^{+}_{i}$ that lies in the plane $z=i$ in step 1
tunnel roof the part of tunnel which is not in the bottom, see Figure 5(c)
left/right roof path tunnel paths shared by roof and bottom.
central roof path tunnel path that is disjoint from left/right roof path
$e_{i}=(s_{i},t_{i})$ special edge of $C$ that is meant to represent $\ell_{i}$
$u$ apex (taking the role of an observer high above)
$H$ graph in the third step of the construction, see black graph in Figure 7
$G^{+}$ an essentially $3$-connected planar graph induced by the neighborhood of $u$
$f_{0}$ outer face of $G^{+}$
$G$ subgraph of $G^{+}$ that is inside the helper triangle
$\overline{C}$ the complex $C$ consists of two copies of $\overline{C}$ with the apex identified
$\varphi$ geometric embedding of $C$
tent Figure 9(b)
tunnel loop $o$ triangle that is shared by two sections
rafter two roof edges of a tunnel loop
wedge $W_{i}$ defined by the apex $u$ and $e_{i}$, Figure 10
plane $P_{i}$ defined by the apex $u$ and $e_{i}$
$p>e_{i}$ segment $pu$ does not cross $e_{i}$
$p<e_{i}$ segment $pu$ does cross $e_{i}$
$S$ sphere around apex $u$
$D$, $D^{+}$ projection of $\varphi(G^{+})$ onto $S$ yielding a drawing of $G$ and $G^{+}$
$a_{i}$ projection of $\varphi(e_{i})$ onto $S$
$\mathcal{A}$ arc arrangement of all $a_{i}$ Figure 11(b)